### v2

# Learning Successor Features with Distributed Hebbian Temporal Memory

Abstract
This paper presents a novel approach to address the challenge of online temporal memory learning for decision-making under uncertainty in non-stationary, partially observable environments. The proposed algorithm, Distributed Hebbian Temporal Memory (DHTM), is based on factor graph formalism and a multicomponent neuron model. DHTM aims to capture sequential data relationships and make cumulative predictions about future observations, forming Successor Features (SF). Inspired by neurophysiological models of the neocortex, the algorithm utilizes distributed representations, sparse transition matrices, and local Hebbian-like learning rules to overcome the instability and slow learning process of traditional temporal memory algorithms like RNN and HMM. Experimental results demonstrate that DHTM outperforms LSTM and a biologically inspired HMM-like algorithm, CSCG, in the case of non-stationary datasets. Our findings suggest that DHTM is a promising approach for addressing the challenges of online sequence learning and planning in dynamic environments.
1 Introduction
Modelling sequential data is one of the essential tasks in Artificial Intelligence as it has many applications, including decision-making and world models (Moerland et al., 2023), natural language processing (Min et al., 2021), conversational AI (Dwivedi et al., 2023), time-series analysis (Eraslan et al., 2019), and video and music generation (Ji et al., 2020). One of the classical approaches to modelling sequential data is forming a representation that stores and condenses the most relevant information about a sequence, and finding a general transformation rule of this information through the dimension of time (Lipton et al., 2015; Harshvardhan et al., 2020; Mathys et al., 2011). We refer to the class of algorithms that use this approach as Temporal Memory (TM) algorithms, as they essentially model the cognitive ability of complex living organisms to remember the experience and make future predictions based on this memory (Hochreiter and Schmidhuber, 1997; Friston et al., 2016, 2018; Parr and Friston, 2017).
This paper addresses the problem of online sequence learning for planning and decision-making under uncertainty, which can be formalized as Reinforcement Learning (RL) for a Partially Observable Markov Decision Process (POMDP) (Poupart, 2005; Singh et al., 2021). Inferring the hidden state in a partially observable environment is, in effect, a sequence modelling problem as it requires processing a sequence of observations to get enough information about hidden states. One of the most efficient representations of the hidden states for discrete POMDP is the Successor Representation (SR) that disentangles hidden states and goals given by the reward function (Dayan, 1993; Samuel J. Gershman, 2018; Millidge and Buckley, 2022). The Successor Features framework is an extension of the SR into continuous POMDP, which employs the same idea of value function decomposition but, instead, for features of a hidden state (Barreto et al., 2017; Touati et al., 2022). Temporal Memory (TM) algorithms can be leveraged to make cumulative predictions about future states and their features, forming SR or SF. This work shows that the proposed algorithm, Distributed Hebbian Temporal Memory (DHTM), can effectively form SFs for navigation tasks in Gridworld and AnimalAI (Fountas et al., 2020) environments.
The most prominent TM algorithms, like a Recurrent Neural Network (RNN) (Boden, 2002) or a Hidden Markov Model (HMM) (Eddy, 2004), use backpropagation to capture data relationships, known for their instability due to recurrent non-linear derivatives. They also require having complete sequences of data at hand during the training. Although the gradient vanishing problem can be partially circumvented in a way, Receptance Weighted Key Value (RWKV) (Peng et al., 2023) or Linear Recurrent Unit (LRU) (Orvieto et al., 2023) models do, the problem of online learning, i.e., learning on non-stationary datasets, is still a viable topic. In contrast to HMM, RNN models and their descendants also lack a probabilistic theory foundation, which is beneficial for modelling sequences captured from stochastic environments (Salaün et al., 2019; Zhao et al., 2020). There is little research on TM models that can be used in fully online adaptable systems interacting with partially observable stochastic environments with access only to one sequence data point at a time, a prevalent case in Reinforcement Learning (Jahromi et al., 2022).
We propose a Distributed Hebbian Temporal Memory (DHTM) algorithm based on the factor graph formalism and multi-compartment neuron model. The resulting graphical structure of our model is similar to one of the Factorial-HMM (Ghahramani and Jordan, 1995) but with a factor graph forming online during training. An important feature of our model is that transition matrices for each factor are stored as different components (segments) of artificial neurons, which makes computations very efficient in the case of sparse transition matrices. Our TM forms sequence representations entirely online and employs only local Hebbian-like learning rules (Hebb, 2005; Churchland and Sejnowski, 1992; Lillicrap et al., 2020), circumventing gradient drawbacks and making the learning process much faster than gradient methods. This comes at the expense of hidden state generalization, as DHTM’s hidden state encodes each observation sequence uniquely. Different trajectories leading to the same position in partially observable Gridworld may result in different hidden states inferred by DHTM. It’s desirable if we want to store different experiences independently, like in human episodic memory (Zhang et al., 2021; W. Schulz and Robins, 2023). Moreover, we argue that generalization is unnecessary when the reward function is well-defined in observation or feature space, and it’s more efficient to use fast hash-like encodings.
The DHTM model notoriously fits Successor Features in the Reinforcement Learning setup. The proposed TM is tested as a world model (Ha and Schmidhuber, 2018; Hafner et al., 2023) for an RL agent architecture, navigating in a Gridworld environment and a more challenging AnimalAI testbed (Crosby et al., 2020). Our algorithm outperforms a classic LSTM and a biologically-inspired HMM-like world model CSCG (George et al., 2021) in online Successor Feature formation task due to a combination of fast Hebbian-like learning and sparse hidden state coding. Another advantage of our algorithm is that it allows its implementation for neuromorphic processors, using only local learning rules.
Our contribution to this work is the following:
  • We propose a distributed memory model (DHTM) based on factor graph formalism and a multicompartment neural model.
  • Our model stores sparse factor functions in neural segments, which significantly lessens the number of trainable parameters, and speeds up learning.
  • The DHTM learns fully online, employing only local Hebbian-like rules.
  • Tested as a world model for an RL agent architecture in Gridworld and AnimalAI, DHTM outperforms LSTM and CSCG in online Successor Features formation for navigation tasks.
2 Background
This section provides basic information about some concepts necessary to follow the paper.
2.1 Reinforcement Learning
This paper considers decision-making in a partially observable environment, usually formalized as a Partially Observable Decision Process (Poupart, 2005). A POMDP is defined as a tuple $\mathcal{M}=(S,A,P,R,O,D,\gamma)$, where $S$—state space, $A$—action space, $P(s,a,s`)=Pr(s`\ |\ s,a)$—transition function, $R(s)$–reward function, O—observation space, $D(a,s`,o)=Pr(o\ |\ a,s`)$—sensor model and $\gamma\in[1,0)$—discount factor, given a transition $s,a\to s`$, where $s\in S$, $a\in A$, $o\in O$. If $S,A,O$ are finite, $P,D$ can be viewed as real valued matrices, otherwise, they are conditional density functions. Here we consider deterministic rewards, which depend only on the current state, i.e. $R(s):S\to\mathbb{R}$.
[…]
2.2 Hidden Markov Model
Partially observable Markov process can be approximated by a Hidden Markov model (HMM) with hidden state space $H$ and observation space $O$. $O$ is the same as in $\mathcal{M}$, but $H$ generally is not equal $S$. Variables $H_{t}$ represent an unobservable (hidden) approximated state of the environment which evolves over time, and observable variables $O_{t}$ represent observations that depend on the same time step state $H_{t}$, and $h_{t},o_{t}$ are corresponding values of this random variables. For the sake of simplicity, we suppose that actions are fully observable and information about them is included into $H_{t}$ variables. For the process of length $T$ with state values $h_{1:T}=(h_{1},\ldots,h_{T})$ and $o_{1:T}=(o_{1},\ldots,o_{T})$, the Markov property yields the following factorization of the generative model:
$p(o_{1:T},h_{1:T})=p(h_{1})\prod_{t=2}^{T}p(h_{t}|h_{t-1})\prod_{t=1}^{T}p(o_{t}|h_{t}).$ (1)
In case of discrete hidden state, a time-independent stochastic transition matrix can be learned with Baum–Welch algorithm (Baum et al., 1970), a variant of Expectation Maximization algorithm. To compute the statistics for the expectation step, it employs the forward-backward algorithm, which is a special case of sum-product algorithm (Kschischang et al., 2001).
2.3 Successor Representation
Successor Representations are such representations of hidden states from which we can linearly infer the state value given the reward function (Dayan, 1993). Here, we assume observation and state spaces are discrete.
$V(h_{t}=i)=\mathrm{E}[\sum^{\infty}_{l=0}\gamma^{l}R_{t+l+1}\ |\ h_{t}=i]=\\ =\sum^{\infty}_{l=0}\gamma^{l}\sum_{j}p(h_{t+l+1}=j\ |\ h_{t}=i)R_{j}=\\ =\sum_{j}\sum^{\infty}_{l=0}\gamma^{l}p(h_{t+l+1}=j\ |\ h_{t}=i)R_{j}=\sum_{j}M_{ij}R_{j},$ (2)
where $\gamma$ is a discount factor, vector $\mathrm{SR}(h=i)=\{M_{ij}\}_{j}$ is a Successor Representation of a state $i$, and $M_{ij}=\sum^{\infty}_{l=0}\gamma^{l}p(h_{t+l+1}=j\ |\ h_{t}=i)$. $R_{j}$ is a reward for observing the state $j$. That is, SR can be computed by a TM that is able to predict future states. TM algorithms effectively predict observations only for a finite time horizon $T$. Therefore, in order to learn SR, a technique similar to TD learning in standard RL may be employed:
$\delta_{ij}$ $=\sum^{T}_{l=0}\gamma^{l}p(h_{t+l+1}=j\ |\ h_{t}=i))+$ (3)
$+\gamma^{T+1}\sum_{k}M_{kj}p(h_{t+T+1}=k\ |\ h_{t}=i)-M_{ij},$
$M_{ij}$ $\leftarrow M_{ij}+\alpha\delta_{ij},$ (4)
where $\alpha\in(0,1)$ is a learning rate, $\delta_{ij}$—TD error for SR.
In partially observable environments, however, exact state values are not known, therefore we operate with state distributions or so-called belief states (Poupart, 2005), which are inferred from observations. In that case, state value and SR are functions of hidden state variable distribution (see details in Appendix B).
2.4 Sparse Distributed Representations
In our work, we design our model to operate with sparse distributed representations (SDRs) to reflect the spatiotemporal property of cortical network activity (Perin et al., 2011). In the discrete time case, SDR is a sparse binary vector in a high-dimensional space. To encode observed dense binary patterns to SDRs, we use a biologically plausible k-WTA (k-winners take all) neural network algorithm called spatial pooler with a Hebbian-like unsupervised learning method (see details in Appendix A).
3 Distributed Hebbian Temporal Memory
3.1 Factor Graph Model
Distributed Hebbian Temporal Memory is based on the sum-product belief propagation algorithm in a factor graph (see Figure 1). Analogously to Factorial-HMM (Ghahramani and Jordan, 1997), we divide the hidden space $H$ into subspaces $H^{k}$. There are four sets of random variables (RV) in the model: $H^{i}_{t-1}$—latent variables representing hidden states from the previous time step (context), $H^{k}_{t}$—latent variables for the current time step, $\Phi^{k}_{t}$—feature variables, and $O^{lm}_{t}$—observable variables. Except for $O^{lm}_{t}$, all random variables have a categorical distribution. In contrast, $O^{lm}_{t}$, are Bernoulli variables because they represent pixels from a binary input image observation. RV state values are denoted as corresponding lowercase letters: $h^{i}_{t-1}$, $h^{k}_{t}$, $\varphi^{k}_{t}$, $o^{lm}_{t}$.
Each variable $\Phi^{k}_{t}$ is considered independent and has a separate graphical model for increased computational efficiency. However, hidden variables of the same time step are statistically interdependent in practice. We introduce their interdependence through a segment computation trick that goes beyond the standard sum-product algorithm (see Eq. 8).
The model also has three types of factors: $M^{i}_{t-1}$—messages from previous time steps, $F^{k}_{c}$—context factor (generalized transition matrix), $F^{k}_{e}$—emission factor. We assume that messages $M^{i}_{t-1}$ include posterior information from the time step $t-1$, therefore we don’t depict observable variables for previous time steps in Figure 1.
Figure 1: Partial factor graph for the DHTM. The input to the model is a sequence of binary images, each pixel is modelled as Bernoulli random variable $O^{lm}_{t}$, where $l$ and $m$ denote corresponding rows and cols of the image. The encoder block forms image categorical features $\Phi^{k}_{t}$ in an unsupervised manner. Each feature $\Phi$ has its own explaining hidden variable, which may depend on hidden variables of the other features and on itself from the previous time step. $F^{k}_{c}$ and $F^{k}_{e}$ are context and emission factors for the corresponding variables. Unary factors $M^{i}_{t-1}$ called messages represent accumulated information about previous time steps.
Further, we discuss only the upper block of the graph in Figure 1, which is DHTM itself. The lower block—an encoder—is described in the Appendix A. The only requirement for the encoder is that its output should be represented as states of categorical variables (features) for the current observation.
3.2 Neural Implementation
The main routine of the DHTM is to estimate distributions of currently hidden state variables given by the equation 5, the computational flow of which is schematically depicted in Figure 2:
$p(h_{t}^{k})\propto\sum_{\Omega_{k}}\prod_{i\in\omega_{k}}M_{t-1}^{i}(h_{t-1}^{i})F_{c}^{k}(h_{t}^{k},\Omega_{k}),$ (5)
where $\Omega_{k}=\{h^{i}_{t-1}:i\in\omega_{k}\}$, $\omega_{k}={i_{1},\dots,i_{n}}$—set of previous time step RV indexes included in $F_{c}^{k}$ factor, $(n+1)$—factor size.
Figure 2: Neuronal implementation of the DHTM. Random variables are represented by cell clusters (white circles), where each cell corresponds to a state and its spike frequency—to the probability of the state $p(h^{k}_{t})$. Cell’s dendritic segments $seg(k)$ correspond to context factor values $f_{l}$ for a particular combination of states (active presynaptic cells) $rec(l)$. Segments’ excitations $E_{l}$ are combined to determine cell’s spike frequency $p(h^{k}_{t})$. Segment’s synaptic weights reflect specificity of $rec(l)$ combination for the segment. Emission factors $F^{k}_{e}$ are fixed and represented by minicolumns inside a variable.
For computational purposes, we translate the problem to the neural network architecture with Hebbian-like learning (for biological interpretation of the model, see Appendix C). As can be seen from Figure 2, every RV can be viewed as a set of spiking neurons representing the RV’s states, that is, $p(h^{k}_{t})=p(c^{j}_{t}=1)$, where $j$—index of a neuron corresponding to the state $h^{k}_{t}$. Cell activity is binary $c^{j}_{t}\in\{0,1\}$ (spike/no-spike), and the probability might be interpreted as a spike rate. Factors $F^{k}_{c}$ and $M^{i}_{t-1}$ can be represented as vectors, where elements are factor values for all possible combinations of RV states included in the factor. Let’s denote elements of the vectors as $f_{l}$ and $m_{u}$ correspondingly, where $l$ corresponds to a particular combination of $k,h^{k}_{t},h^{i_{1}}_{t-1},\dots,h^{i_{n_{l}}}_{t-1}$ state values and $u$ indexes all neurons representing states of previous time step RVs.
Drawing inspiration from biological neural networks, we group a neuron’s connections into dendritic segments. A segment acts as an independent computational unit that detects a particular input pattern (a context state) defined by its own receptive field. In our model, a segment links together factor value $f_{l}$, the computational graph shown in Figure 2, and the excitation $E_{l}$ induced by the segment $l$ to the cell it is attached to. The segment is active, i.e., $s_{l}=1$ if all its presynaptic cells are active; otherwise, $s_{l}=0$. Computationally, a segment transmits its factor value $f_{l}$ to a cell it is attached to if the context matches the corresponding state combination.
We can now rewrite equation 5 as the following:
$p(h_{t}^{k})\propto\sum_{l\in\mathrm{seg}(j)}L_{l}f^{k}_{l},$ (6)
where $L_{l}=\prod_{u\in\mathrm{rec}(l)}m_{u}$ is segment’s likelihood as long as messages are normalized to probability distributions, $\mathrm{seg}(j)$—indexes of segments that are attached to cell $j$, $\mathrm{rec}(l)$—indexes of presynaptic cells that constitute receptive field of a segment with index $l$.
Initially, all factor entries are zero, meaning cells have no segments. As learning proceeds, new non-zero connections grouped into segments are grown. In equation 6 we benefit from having sparse factor value vectors because its complexity depends linearly on the amount of non-zero components. And that’s usually the case in our model due to one-step Monte-Carlo learning and specific form of emission factors $F^{k}_{e}$:
$F^{k}_{e}(h^{k}_{t},o^{k}_{t})=\mathbb{I}[h^{k}_{t}\in\mathrm{col}(\varphi^{k}_{t})],$ (7)
where $\mathbb{I}$—indicator function, $\mathrm{col}(\varphi^{k}_{t})$ is a set of hidden states connected to the feature state $\varphi^{k}_{t}$ that forms a column. The form of emission factor is inspired by presumably columnar structure of the neocortex and was shown to induce sparse transition matrix in HMM (George et al., 2021).
Segment likelihood $L_{l}$, resulting from the sum-product algorithm, is calculated as if presynaptic cells are independent. However, it’s not usually the case for sparse factors. To take into account, approximately, their interdependence, we substitute the following equation for segment log-likelihood:
$\log L_{l}$ $=w_{l}\log\frac{1}{n_{l}}\sum_{u\in\mathrm{rec}(l)}w_{ul}m_{u}+$
$+\sum_{u\in\mathrm{rec}(l)}(1-w_{ul})\log m_{u},$ (8)
where $w_{pl}$—synapse efficacy or neuron specificity for segment, such that $w_{ul}=p(s_{l}=1|c^{u}_{t-1}=1)$, $w_{l}=\frac{1}{n_{l}}\sum_{u}w_{ul}$—average synapse efficacy for segment $l$, and $n_{l}$-number of cells in segment’s receptive field.
The idea that underlies the formula is to approximate between two extreme cases:
  • $p(s_{l}=1|c^{u}_{t-1}=1)\to 1$ for all $u$, which means that all cells in the receptive field are dependent and are part of one cluster, i.e., they fire together. In that case, it should be $p(s_{l})=m_{u}$ for any $u$, but we also reduce prediction variance by averaging between different $u$.
  • $p(s_{l}=1|c^{u}_{t-1}=1)\to 0$ for all $u$ means that presynaptic cells don’t form a cluster. In that case, segment activation probability is just a product of the activation probability of each cell.
The resulting equation for belief propagation in DHTM is the following:
$p(h^{k}_{t})=p(c^{j}_{t}=1)=\underset{j\in\mathrm{cells}[H^{k}_{t}]}{\mathrm{softmax}}(\underset{l\in seg(j)}{\mathrm{max}}(E_{l})),$ (9)
where $E_{l}=\log f_{l}+\log L_{l}$, $\mathrm{cells}[H^{k}_{t}]$—indexes of cells that represent states for $H^{k}_{t}$ variable. Here, we also approximate logarithmic sum with $\mathrm{max}$ operation inspired by the neurophysiological model of segment aggregation by cell (Stuart and Spruston, 2015).
The next step after computing $p(h^{k}_{t})$ distribution parameters is to incorporate information about current observations $p(h^{k}_{t}\ |\ o^{k}_{t})\propto p(h^{k}_{t})\mathbb{I}[h^{k}_{t}\in\mathrm{col}(o^{k}_{t})]$. After that, the learning step is performed. The step for closing the loop of our TM algorithm is to assign the posterior for the current step $p(h^{k}_{t}\ |\ o^{k}_{t})$ to $M^{i}_{t-1}$.
DHTM learns $f_{l}$ and $w_{ul}$ weights by Monte-Carlo Hebbian-like updates. First, $h^{i}_{t-1}$ and $h^{k}_{t}$ are sampled from their posterior distributions: $p(h^{i}_{t-1}\ |\ o^{i}_{t-1})\propto M^{i}_{t-1}$ and $p(h^{k}_{t}\ |\ o^{k}_{t})$ correspondingly. Then $f_{l}$ is updated according to the segment’s $s_{l}$ and its cell’s $c^{j}_{t}$ activity so that $f_{l}$ is proportional to several coincidences $s_{l}=c^{j}_{t}=1$ during the recent past, i.e., cell and its segment are active at the same time step:
$\Delta f_{l}=\alpha(c^{j}_{t}-f_{l})s_{l},$ (10)
where $\alpha\in[0,1)$ is segment’s learning rate. That is, $f_{l}$ is trained in that way that it is equal to the average cell’s activation frequency, given the segment is active. It’s similar to Baum-Welch’s update rule (Baum et al., 1970) for the transition matrix in HMM, which, in effect, counts transitions from one state to another, but, in our case, the previous state (context) is represented by a group of RVs, not just one hidden RV.
Weights $w_{ul}$ are also updated by the Hebbian rule to reflect the specificity of a presynaptic $u$ for activating a segment $l$. That is, they are targeted to represent probability $p(s_{l}=1\ |\ c^{u}_{t-1}=1)$ that segment $s_{l}$ is active, given cell $u$ was active at the previous time-step. We could learn it by counting activation coincidences and mismatches. But in our algorithm it is approximated as exponential moving average of segment’s $s_{l}$ frequency activation, given $c^{u}_{t-1}=1$: $\Delta w_{ul}=\beta\cdot\mathbb{I}[c^{u}_{t-1}=1]\cdot(\mathbb{I}[s_{l}=1]-w_{ul})$, where $\beta\in[0,1)$ — learning rate.
3.3 Agent Architecture
We incorporate DHTM as a part of an RL agent. We use the same agent for other memories tested as well. The agent consists of a memory model and a feature reward function. The memory model generates SF by predicting cumulative future distributions of feature variables $\Phi^{k}$ according to equation 12. Here, we assume that feature representations of observations are good enough to define the reward function $r(\varphi^{k}_{t})$, which is also learned during interaction with the environment and, combined with SF representations, is used to estimate the action value function. For more details on employed value function decomposition, see Appendix B.
Algorithm 1 General agent training procedure
1:  for episode=1..n do
2:     RESET_MEMORY()
3:     action $\leftarrow$ initial_action
4:     while (not terminal) (steps $<$ max_steps) do
5:        obs, reward $\leftarrow$ STEP()
6:        features $\leftarrow$ ENCODE(PREPROCESS(obs))
7:        OBSERVE(features, action)
8:        REINFORCE(reward, features)
9:        action $\leftarrow$ SAMPLE_ACTION()
10:        ACT(action)
11:     end while
12:  end for
The agent training procedure is outlined in Algorithm 1. For each episode, the memory state is reset to a fixed initial message with RESET_MEMORY() and action variable is set to a fixed initial action. An observation image returned by an environment (obs) is first preprocessed to get events, mimicking a simple event-based camera with a floating threshold determined from the average difference between the current and previous step image intensities. The resulting events are encoded to SDRs with a biologically inspired spatial pooling encoder described in Appendix A. In OBSERVE() routine, the memory learns to predict next feature states as described in Section 3. An agent learns associations to feature states and rewards in REINFORCE() function:
$r^{k}_{i}\leftarrow r^{k}_{i}+\alpha\mathbb{I}[\varphi^{k}_{t}=i](R_{t}-r^{k}_{i})$ (11)
where $\alpha$ is a learning rate, $R_{t}$—a reward for the current time step.
We include actions in the model by forcing some of the hidden variables $H^{k}_{t}$ to represent actions. That is, we assume that information about action is included in the hidden state of the model. For example, if we have 4 actions, we set 4 states for one of the hidden variables and set its state from observation of the action.
In SAMPLE_ACION(), for each action and current belief state, TM predicts the next belief states for which SFs of horizon $T$ are formed using uniform strategy, i.e. setting action variable to uniform distribution. In most experiments, $T$ is dynamically determined from $\Phi^{k}_{t+l}$ distribution. If the distribution of $\Phi^{k}_{t+l}$ is close to uniform or states with high positive reward have high probability, the prediction cycle is stopped, rendering the prediction process similar to breadth-first search.
$\mathrm{SF}_{t+T}(\varphi^{k}=j\ |\ o_{0:t},a_{t})=\\ =\sum^{T}_{l=0}\gamma^{l}\sum_{i}p(\varphi^{k}_{t+l+1}=j\ |\ h^{k}_{t}=i)p(h^{k}_{t}=i\ |\ o_{0:t},a_{t})$ (12)
$Q(o_{0:t},a_{t})=\sum_{jk}\mathrm{SF}_{t+T}(\varphi^{k}=j\ |\ o_{0:t},a_{t})r^{k}_{j},$ (13)
where $o_{0:t}$—observation history.
4 Experiments
We test our model in a reinforcement learning task in Gridworld environment, where successor features are easy to interpret and the encoder block (see Sec. 3) is not required, and in a more challenging AnimalAI 3D environment. This section shows how different memory models affect SF learning and an RL agent’s adaptability. In our work, we compare the proposed DHTM model with LSTM (Hochreiter and Schmidhuber, 1997) and CSCG (George et al., 2021). For each setup, the results are averaged over at least five experiments with different seed values and standard deviation is shown. See Appendix D for details about baselines, their model parameters and training regimes. In Appendix E, environments and setups are more thoroughly described. The source code of experiments and algorithms’ implementations is available at https://anonymous.4open.science/r/dhtm-D02F.
4.1 Gridworld
First we test DHTM in MDP setup, that is, observation state corresponds to position in 5x5 maze. Each trial consists of two phases. In the first phase, lasting 500 episodes, there are no obstacles on the way to the terminal state that gives a reward to the agent. In the second phase, the terminal state is surrounded by a wall, so that the agent has to detour in order to get the reward, which remains in the same position. Each episode starts with the agent in the same position and ends when terminal state is entered (i.e. the reward is collected) or maximum action steps reached. Each episode, we measure the amount of action steps for agent required to reach the goal.
The results show (Fig. 3) that the agent that uses DHTM for SR formation faster converges to the optimal trajectory than CSCG. CSCG is able to converge to optimal trajectory only if the initial training dataset is gathered using random strategy after environmental changes. It requires an oracle that tells when changes occurred to gather a new training dataset. Otherwise, the agent with CSCG memory converges to suboptimal trajectories as the training batch gets very correlated trajectories following the current agent’s policy, which causes overfitting. Here, we also provide results of Q-table and SR-table agents for reference.
Figure 3: Amount of steps required for the agent to reach the goal in fully observable 5x5 Gridworld using different types of memory: DHTM and CSCG. CSCG is trained in two modes. The first mode is with oracle, which tells the agent to gather some trajectories using uniform policy after environment changes. The second mode is without oracle, so the agent continues using its current strategy to gather new trajectories. At the 500th episode, the terminal state is blocked by a wall, so the agent needs to find a new way to the goal. There are also plots of learning curves for Q-table and SR-table agents for reference.
The second test was conducted in partially observable Gridworld in a 5x5 changing maze with a stationary terminal state as before. However, now all the agent can see is the colour of the cell in which it is located. The agent is also able to see the colour of an obstacle it encounters, receiving additional punishment.
As can be seen from Figure 4, in contrast to DHTM, CSCG and LSTM are not able to converge after reward is blocked at the 300th episode. Even with an oracle signalling about changes in the environment, which enables to gather new fully random trajectories, neither CSCG nor LSTM are able to relearn maze’s structure in online case scenario. It can be seen, however, that DHTM converges to a suboptimal trajectory of length 13 on average vs 9 steps for optimal trajectory, after reward is blocked. This may be the result of inferior generalisation properties of DHTM, which we discuss in Section 5.
Figure 4: Learning curves measured in action steps until the goal for the agent using different memory types for SF formation in partially observable 5x5 Gridworld environment. In the 300th episode, the goal is blocked by a wall, so the agent needs to find a new optimal trajectory. CSCG is trained in two modes. The first mode is with an oracle, which tells the agent to gather some trajectories using uniform policy after the environment changes. The second mode is without an oracle, so the agent continues using its current strategy to gather new trajectories.
In these experiments, we aimed to show that in cases of quite limited and non-stationary trajectory buffer size, an RL agent armed with DHTM memory is more adaptive than those using LSTM and CSCG. However, we do not claim that it is impossible, in principle, to train LSTM and CSCG in such a setup. We only show that DHTM notoriously fits this kind of task without numerous workarounds that would be required by those classical methods.
4.2 AnimalAI
AnimalAI is a 3D environment that is able to give an agent rich visual input in the form of RGB images. The objective of the experiment described here is to give a hint that DHTM is also able to deal with complex visual input due to its encoder block described in Appendix A.
The agent’s task is to reach a ball, which is food, that is located in the same room, which is 10 by 10 meters. Start agent and food positions are the same between episodes. All the agent needs to do is find the shortest path to the goal, as in the Gridworld experiments. In each episode, we measure the number of actions (steps) the agent takes to reach the goal.
The results are shown in Figure 5. We can see that after 150 episodes with the maximum number of steps set to 40, the agent stably reaches the goal. Although slight instability persists, the agent’s path is near optimal.
We weren’t able to train CSCG and LSTM with the same encoder block as we used for DHTM, even after increasing their trajectory buffer sizes. We think that the problem might be because of the non-stationarity of the encoder’s output, which adapts continuously.
Figure 5: Amount of action steps the agent uses to reach the goal in a 10x10-meter room in AnimalAI environment using DHTM memory for SF formation. The agent and the goal have the same initial positions throughout the episodes. Each timestep, the agent gets an image of a first-person view of the environment and produces one of three actions: turn left, turn right, or go forward.
5 Conclusion
This paper introduces a novel probabilistic Factorial-HMM-like algorithm, DHTM, for learning an observation sequence model in non-stationary environments that uses local Hebbian-like learning rules. DHTM is scalable to multiple feature variables by employing sparse distributed representations and sparse factor function implementation using segments, which inspire biologically plausible multicomponent neural models. In contrast to methods that use Monte-Carlo trajectory sampling for future state probability estimation, our method can perform belief propagation, so each prediction step adds a constant amount of computation, allowing a long planning horizon.
Experiments show that our memory model can quickly learn and unlearn observation sequences and transition dynamics, even in changing environments. In that case, methods that rely on backpropagation, like LSTM and CSCG, are difficult to train and require rigorous trajectory buffer management, a common problem in RL tasks.
One of the main limitations of the DHTM is that its temporal context is completely random, as it is formed on the fly using only the local information available. As a result, the mechanism of context formation allows only limited generalizations between different trajectories. However, this is what makes fast learning possible. Combining this fast learning with more slow learning that allows generalizations is a promising direction. Forming Successor Features combined with a two-level hierarchy of DHTM layers may provide the next step to a more generalized sequence model. The first (fast) DHTM layer may be used to form Successor Features both for planning and as current state representation, which is then observed by the second (slow) DHTM layer. We hypothesize that because SFs should provide a better generalization than the raw features, the second layer should be able to form a more versatile model of the environment.
6 Impact Statements
This paper presents research that aims to advance the field of machine learning. There are numerous potential societal consequences of our work, none of which we believe require special emphasis here.
Appendix A Encoding and Decoding Observations
Because our model is designed to work with sparse distributed representations and the testing environments do not provide observations as SDRs by default, an encoding procedure is required. For this task, we use a modified version of the Spatial Pooler (SP) [Cui et al., 2017, Mnatzaganian et al., 2017], a distributed noise-tolerant online clustering neural network algorithm that converts input binary patterns into SDRs with fixed sparsity while retaining pairwise similarity [Kuderov et al., 2023]. The SP algorithm learns a spatial specialization of neurons’ receptive fields using the local Hebbian rule and k-WTA ($k$ winners take all) inhibition [Oster et al., 2009]. Here we outline the main differences from the “vanilla” version of the SP algorithm described in Cui et al. [2017].
During an agent’s decision-making process pipeline, the SP encoder accepts a current observation $o$ and transforms it to a latent state SDR $z$. Observation $o$ should be an array of binary states. To transform an RGB image into a binary array, we perform the following processing steps, simulating a simple event-based camera:
  1. 1.
    Transform image to greyscale.
  2. 2.
    Calculate difference between current and previous greyscale image $d$.
  3. 3.
    Determine average difference over the entire image $\tau=\mathrm{mean}(d)$.
  4. 4.
    For pixels that have $d$ above $\tau$ set $1$, for others set 00.
In terms of processing, our SP encoder functions as a standard artificial neural network with a k-WTA binary activation function.:
$\text{overlaps}_{i}$ $=\beta_{i}W_{i}o$ (14)
$z_{i}$ $=\mathbb{I}\left[i\in\text{kWTA}(\text{overlaps})\right],$ (15)
where $o$—a binary observation vector, $W_{i}$—a row-vector representing $i$-th neuron’s connection weights (where non-existing connections have zero weights), $\text{overlaps}_{i}$—a value representing the strength of the input pattern recognition with the neuron $i$  [^1] While the name “overlap” does not exactly reflect its meaning in our SP modification, because it is not a binary overlap between a receptive field and an input pattern, we kept it on purpose to refer to the similar term commonly used for the original SP., $\beta_{i}$—an $i$-th neuron boosting value, $z_{i}$—an $i$-th bit of an output SDR, $\mathbb{I}[\dots]$—an indicator function, kWTA—a $k$-winners-take-all activation function returning $k$ indices of the neurons with the highest overlap.
One difference between the “vanilla” SP algorithm and ours is that we do not distinguish between potential and active neural connections. Because all [existing] connections are active, they all participate in calculating overlaps. In the overlaps calculation, non-binary, that is, real-valued weights are used, similar to artificial neural networks, as shown in equation 14. Furthermore, each neuron has a fixed capacity to produce neurotransmitters, which it distributes between its synaptic connections. This means that we keep all neuron weights normalized and summed to one. While it achieves the same Hebbian learning with homeostatic plasticity as the original SP, the exact formula is slightly different:
$\tilde{W}_{i}$ $=W_{i}+\alpha z_{i}\frac{\text{RF}_{i}\odot o}{\sum_{j}{\text{RF}_{i}\odot o}}$
$W_{i}$ $\leftarrow\frac{\tilde{W}_{i}}{\sum_{j}{\tilde{W}_{j}}},$ (16)
where $\tilde{W}_{i}$—a row of new $i$-th neuron weights before normalization, $\alpha$—learning rate, $z_{i}$—a binary value representing the current activity state of the $i$-th neuron, $RF_{i}$—an $i$-th row of the binary connectivity matrix representing an $i$-th neuron receptive field, $\odot$—elementwise product, $o$—a binary observation vector.
The original SP algorithm has several drawbacks, including encoding instability caused by an innate homeostatic plasticity mechanism known as boosting, which helps neurons specialize and increases overall adaptability but makes memorization tasks more difficult, and slow processing on large inputs such as images, where an encoding overhead becomes noticeable when compared to overall model timings around 1k input size.
The introduction of the newborn stage, which follows the ideas proposed in Dobric et al. [2022], solves an encoding instability problem. The newborn stage of a spatial pooler occurs during the early stages of its learning process, when its neurons are expected to specialize. The boosting, which is intended to aid in the specialization process, is activated only during the newborn stage and its scale gradually decreases from the configured value to zero. Boosting remains turned off during an encoder’s “adulthood”, reducing the possibility of spontaneous re-specialization.
To reduce processing overhead, we use a much more sparsified connection matrix than in the original SP version. We randomly initialize connections with 40-60% sparsity, which is typical for the “vanilla” SP. Then, during the newborn stage, we gradually prune the vast majority of the weakest connections, resulting in neurons that are highly specialized due to their small receptive fields. We typically configure the final receptive field size in relation to the average input pattern size (usually 25-200% of it, resulting in 0.1-10% connections sparsity). For example, if binary input patterns have on average 100 active bits out of 1000, we can set the target size of receptive fields to 25, which is 25% of the active input size and corresponds to 2.5% connection matrix sparsity. As a result, the spatial pooler’s instability (and thus adaptiveness!) becomes even more limited in the adult stage.
Because of its soft discretization (from the distributed representation) and clusterization properties, we expect SP to assist the model with input sequence memorization and an environment transition dynamics generalization tasks in addition to the encoding itself. However, because the SP encoder learns online, particularly during the newborn stage, its output representation can be highly unstable during the early stages, potentially resulting in a performance drop.
To visualize and debug an encoded observation, we also learn a decoder, which is a linear neural layer learned locally with gradient descend on the MSE error between the predicted reconstruction and the actual observation.
Appendix B Value Function Decomposition
In our agent model, we approximate the reward function R(s) as a sum: $R_{t}=\frac{1}{n}\sum^{n}_{k=1}{r(\varphi^{k}_{t})\varphi^{k}_{t}}$, where $r(\varphi^{k}_{t})$ is a reward associated with state $\varphi^{k}_{t}$, $n$–number of feature variables. Then, similarly to the Successor Representation idea (see Section 2.3), the value function can be represented as:
$V(h_{t})$ $=\mathrm{E}[\sum^{\infty}_{l=0}\gamma^{l}R_{t+l+1}\ |\ h_{t}]=\sum^{\infty}_{l=0}\gamma^{l}\mathrm{E}[\frac{1}{n}\sum^{n}_{k=1}{r(\varphi^{k}_{t})}\ |\ h_{t}]$
$=\frac{1}{n}\sum^{n}_{k=1}\sum_{j}\sum^{\infty}_{l=0}\gamma^{l}p(\varphi^{k}_{t+l+1}=j\ |\ h^{k}_{t})r^{k}_{j}$
$=\frac{1}{n}\sum^{n}_{k=1}\sum_{j}M^{k}_{j}(h^{k}_{t})r^{k}_{j},$ (17)
where $M^{k}_{j}(h^{k}_{t})=\sum^{\infty}_{l=0}\gamma^{l}p(\varphi^{k}_{t+l+1}=j\ |\ h^{k}_{t})$, $h_{t}=(h^{1}_{t},...,h^{n}_{t})$—hidden state vector of variables $\{H^{k}_{t}\}_{k}$.
Appendix C Biological Interpretation
Neural implementation of the DHTM is inspired by neocortical neural networks (see Fig. 6). Hidden variables $H^{k}$ may be considered as populations of excitatory pyramidal neurons in cortical layer L2/3 of somatosensory areas, with lateral inhibition modelled as $\mathrm{softmax}$ function. Staiger and Petersen [2021] showed that neurons in this layer are responsible for temporal context formation.
The neuronal activity at timestep $t$ can be thought to carry messages $M^{k}_{t-1}$. Messages are propagated through synapses of dendritic segments, which correspond to factors $F^{k}_{c}$. Dendritic segments of biological neurons are known to be coincidence detectors of its synaptic input [Stuart and Spruston, 2015]. We use the notion of dendritic segment to sparsely represent context factors $F^{k}_{c}$, as each factor value corresponds to a particular combination of states (or active cells).
Feature variables $\Phi^{k}_{t}$ may be considered to represent cells of a granular layer (L4), as they are known to be the main hub for sensory excitation for L2/3. L2/3 cells that have common sensory input from the layer L4 are modelled as columns for particular feature states $\mathrm{col(\varphi^{k}_{t})}$ [Mountcastle, 1997].
Figure 6: Biological view of the neural implementation of the DHTM. Variables $H^{\cdot}_{t-1}$ correspond to populations of neurons that have common sensory input and lateral inhibitory competition. Dendritic segments correspond to factor values $f_{l}$. spike frequency of a neuron reflects state probability $p(h^{k}_{t})$ of a variable.
Appendix D Implementation details
As mentioned in Section 3.3, we incorporate DHTM as a part of an RL agent, which has a memory model and a feature reward function. All tested memory models share the same pipeline—they learn sequences of encoded binary observations (i.e. SDRs that we get from Spatial Pooler encoder described in Appendix A) concatenated with one-hot encoded actions.
D.1 LSTM
LSTM baseline was implemented with a single LSTMCell from PyTorch library [Paszke et al., 2019]. It is supported by an additional symexp-layer to encode input before passing it to the LSTM cell and a symexp-layer to decode the LSTM cell’s output from the LSTM’s hidden state back to the input representation, where symexp activation function, $\mathrm{symexp}(x)=\mathrm{sign}(x)e^{\lvert x\rvert-1}$, is a reverse of symlog function: $\mathrm{symlog}=\mathrm{sign}(x)\log{(\lvert x\rvert+1})$.
For the Gridworld task, LSTM with a hidden state size of 100 is trained on a buffer of 100 trajectories with a decreasing learning rate starting at 0.02. 25 trajectories are left out for validation on each epoch. The training is terminated when the validation loss is not decreasing more than 3 epochs in a row. After that, 30% of the buffer is replaced with new trajectories, and training continues.
We also incorporated some notion of random variables in hidden states by splitting the hidden state of the tested LSTM into groups. In all experiments the hidden state represents $10$ categorical variables with $10$ states. That is, LSTM is forced to learn $10$ categorical distributions with multi-cross-entropy loss to explain the observed sequences, which is a somewhat close to the multi-categorical hidden state representation used in DreamerV2/V3 [Hafner et al., 2023]. The idea of using symexp activation function, mentioned above, is inspired by Dreamer too, and is used to remedy the problem of learning extreme probability values. Without symexp the neural network has to represent zero probability with high negative logit values and one-probability with high positive logit values, which is hard to reach with low learning rate and may lead to instabilities. Thus, symexp function makes it faster to reach target values in log space.
D.2 CSCG
CSCG baseline was implemented using code from the repository accompanying the paper (https://github.com/vicariousinc/naturecomm_cscg). CSCG with 10 clones per observation state was trained on batches with a size of 2500 observation steps for Gridworld tasks with 1000 EM algorithm iterations. We iteratively calculated the exponential moving average of transition matrices obtained for different batches with a smoothing coefficient $\alpha=0.9$ as described in the paper [Dedieu et al., 2019]. This smoothed transition matrix was used as an initialization for the next training batch and for inference. The first batch of observations is gathered using the uniform policy, and later batches are gathered using the current best policy. In so-called oracle mode, after the environment changes, the agent gathers a uniform policy batch again.
D.3 DHTM
In the POMDP Gridworld task, a graph for DHTM consists of three hidden variables $H$ with 40 cells per feature state, which is raw observation with 7 states in the case of Gridworld. Each hidden variable gets the same observation, and their predictions are averaged. Using several hidden variables here is essential, as it significantly increases the amount of trajectories memory is able to store by decreasing collision probability for hidden state representations of different trajectories. The same graph is used for the AnimalAI task, but with the encoder block forming one feature variable with 50 states.
In MDP Gridworld, a graph with one hidden variable and one cell for observation state is used, as there is no need to distinguish observations in different contexts.
Generally, the softmax strategy much better fits DHTM than Epsilon greedy. This is connected with the hidden state generalization problem discussed in Section 5. Each new trajectory creates a new hidden state, for which possible future trajectories are not known. The more exploration there is, the less useful the memory is. Softmax renders exploration more narrowed around already-known good trajectories, allowing better reuse of DHTM’s knowledge.
SFs are formed using $\gamma=0.8$ in all experiments except MDP Gridworld, where $\gamma=0.99$ is used. The maximum planning horizon is set to 50 in POMDP Gridworld and AnimalAI, with stopping criteria based on predicted feature variable distribution: 1) DKL between prediction and uniform distribution is less than 0.1, 2) net probability of feature states with positive rewards is greater than 0.01. For MDP Gridworld, planning steps are fixed and equal to 10. The same SF parameters were used for CSCG and LSTM.
In all experiments, segments are initialized with the maximum value, which is $1$. Synapses are initialized with $0.5$. Segments’ and synapses’ learning rates are $0.1$. We also prune segments with values less than $0.3$ to speed up learning.
Appendix E Experimental Setups
E.1 Gridworld
We use a simple homebrew Gridworld implementation. Each Gridworld position maps to a colour, reward, and terminal state indicator. There are two special colours for the world border and obstacles. Each timestep, an agent gets the colour index of the current agent’s position. In MDP mode, the agent gets a tuple of its coordinates in the maze. An agent is able to see obstacles or border colour when trying to go through them. We also punish an agent trying to go through a wall or world border in addition to base step reward punishment to facilitate faster convergence to optimal trajectories.
Setups used in our experiments are depicted in Figures 7 and 8. Here, the red colour corresponds to obstacles and the light gray colour to the reward. In each episode, the agent and the terminal rewarding state have the same initial positions. The episode ends when the terminal state is entered or the limit of action steps is reached.
Figure 7: Schematic view of MDP Gridworld setups, in which agent gets its Gridworld position instead of colour index. The left picture shows the initial setup of the environment, with the reward in the center. The right picture represents the environment after the reward is blocked by obstacles (red squares). Agent’s initial position in the lower left corner.
Figure 8: Schematic view of POMDP Gridworld setups. The left picture shows the initial setup of the environment, with the reward delivered by the light gray square. The right picture represents the environment after the reward is blocked by obstacles (red squares). Agent’s initial position in the lower left corner.
E.2 AnimalAI
AnimalAI is a testbed inspired by experiments with animals [Crosby et al., 2020]. The environment consists of 3D area surrounded by a wall and many different objects that can be placed using a configuration file, including walls, food, ramps, trees, movable obstacles, and more. We tested our agent in a 10x10-meter room surrounded by walls (see Fig. 9). Each timestep, the agent gets an RGB image of a first-person view of the environment and is able to influence its behaviour by choosing one of three actions: turn left, turn right, or go forward. The agent and food item have the same fixed initial positions for each episode. The episode ends if the food item is collected or the maximum episode time is reached. To make an agent’s actions more discretized, we repeat its choice three times and skip three frames.
Figure 9: Sample of the agent’s first-person view of our setup in AnimalAI environment. The green ball represents a goal, a food item to be collected.
Appendix F Glossary
Categorical Random Variable—a discrete random variable that can take on of finite $K$ possible states.
Cortical Column or Minicolumn—a population of neurons in the neocortex that spans across layers and shares sensory input.
Dendritic segment—a group of synapses (neuron’s connections) that acts as an independent computational unit affecting the resulting neuron’s activity.
Factor Graph—bipartite graph representing the factorization of a probability distribution, with one part representing factor nodes and another—random variables.
Hidden Markov Model (HMM)—statistical model of a stochastic process where state probability depends only on previous state of the process.
Multi-compartment neuron model—a model of neuron that divides neuron’s connections into groups (segments) of different types (compartments), where each group may be considered as partly independent computational unit and groups of each compartment may affect the neuron’s activity differently.
Sparse Distributed Representations (SDR)—sparse binary vector in a high-dimensional space, usually formed by k-WTA algorithms.
Spatial Pooler (SP)—a distributed noise-tolerant online clustering neural network algorithm that converts input binary patterns into SDRs with fixed sparsity while retaining pairwise similarity.
Successor Representations (SR)—a discounted sum of future [one-hot encoded] observations.
Successor Features (SF)—a generalization of SR, a discounted sum of future latent states.
Temporal Memory (TM)—in this work by this term we mean “memory for sequences”.

### v3

@@
 1 Introduction
-Modelling sequential data is one of the essential tasks in Artificial Intelligence as it has many applications, including decision-making and world models (Moerland et al., 2023), natural language processing (Min et al., 2021), conversational AI (Dwivedi et al., 2023), time-series analysis (Eraslan et al., 2019), and video and music generation (Ji et al., 2020).
-One of the classical approaches to modelling sequential data is forming a representation that stores and condenses the most relevant information about a sequence, and finding a general transformation rule of this information through the dimension of time (Lipton et al., 2015; Harshvardhan et al., 2020; Mathys et al., 2011).
-We refer to the class of algorithms that use this approach as Temporal Memory (TM) algorithms, as they essentially model the cognitive ability of complex living organisms to remember the experience and make future predictions based on this memory (Hochreiter and Schmidhuber, 1997; Friston et al., 2016, 2018; Parr and Friston, 2017).
-This paper addresses the problem of online sequence learning for planning and decision-making under uncertainty, which can be formalized as Reinforcement Learning (RL) for a Partially Observable Markov Decision Process (POMDP) (Poupart, 2005; Singh et al., 2021).
-Inferring the hidden state in a partially observable environment is, in effect, a sequence modelling problem as it requires processing a sequence of observations to get enough information about hidden states.
-One of the most efficient representations of the hidden states for discrete POMDP is the Successor Representation (SR) that disentangles hidden states and goals given by the reward function (Dayan, 1993; Samuel J.
-Gershman, 2018; Millidge and Buckley, 2022).
-The Successor Features framework is an extension of the SR into continuous POMDP, which employs the same idea of value function decomposition but, instead, for features of a hidden state (Barreto et al., 2017; Touati et al., 2022).
-Temporal Memory (TM) algorithms can be leveraged to make cumulative predictions about future states and their features, forming SR or SF.
-This work shows that the proposed algorithm, Distributed Hebbian Temporal Memory (DHTM), can effectively form SFs for navigation tasks in Gridworld and AnimalAI (Fountas et al., 2020) environments.
-The most prominent TM algorithms, like a Recurrent Neural Network (RNN) (Boden, 2002) or a Hidden Markov Model (HMM) (Eddy, 2004), use backpropagation to capture data relationships, known for their instability due to recurrent non-linear derivatives.
-They also require having complete sequences of data at hand during the training.
-Although the gradient vanishing problem can be partially circumvented in a way, Receptance Weighted Key Value (RWKV) (Peng et al., 2023) or Linear Recurrent Unit (LRU) (Orvieto et al., 2023) models do, the problem of online learning, i.e., learning on non-stationary datasets, is still a viable topic.
-In contrast to HMM, RNN models and their descendants also lack a probabilistic theory foundation, which is beneficial for modelling sequences captured from stochastic environments (Salaün et al., 2019; Zhao et al., 2020).
-There is little research on TM models that can be used in fully online adaptable systems interacting with partially observable stochastic environments with access only to one sequence data point at a time, a prevalent case in Reinforcement Learning (Jahromi et al., 2022).
-We propose a Distributed Hebbian Temporal Memory (DHTM) algorithm based on the factor graph formalism and multi-compartment neuron model.
-The resulting graphical structure of our model is similar to one of the Factorial-HMM (Ghahramani and Jordan, 1995) but with a factor graph forming online during training.
-An important feature of our model is that transition matrices for each factor are stored as different components (segments) of artificial neurons, which makes computations very efficient in the case of sparse transition matrices.
-Our TM forms sequence representations entirely online and employs only local Hebbian-like learning rules (Hebb, 2005; Churchland and Sejnowski, 1992; Lillicrap et al., 2020), circumventing gradient drawbacks and making the learning process much faster than gradient methods.
-This comes at the expense of hidden state generalization, as DHTM’s hidden state encodes each observation sequence uniquely.
-Different trajectories leading to the same position in partially observable Gridworld may result in different hidden states inferred by DHTM.
-It’s desirable if we want to store different experiences independently, like in human episodic memory (Zhang et al., 2021; W.
-Schulz and Robins, 2023).
-Moreover, we argue that generalization is unnecessary when the reward function is well-defined in observation or feature space, and it’s more efficient to use fast hash-like encodings.
-The DHTM model notoriously fits Successor Features in the Reinforcement Learning setup.
-The proposed TM is tested as a world model (Ha and Schmidhuber, 2018; Hafner et al., 2023) for an RL agent architecture, navigating in a Gridworld environment and a more challenging AnimalAI testbed (Crosby et al., 2020).
-Our algorithm outperforms a classic LSTM and a biologically-inspired HMM-like world model CSCG (George et al., 2021) in online Successor Feature formation task due to a combination of fast Hebbian-like learning and sparse hidden state coding.
-Another advantage of our algorithm is that it allows its implementation for neuromorphic processors, using only local learning rules.
-Our contribution to this work is the following:
-  • We propose a distributed memory model (DHTM) based on factor graph formalism and a multicompartment neural model.
-  • Our model stores sparse factor functions in neural segments, which significantly lessens the number of trainable parameters, and speeds up learning.
-  • The DHTM learns fully online, employing only local Hebbian-like rules.
-  • Tested as a world model for an RL agent architecture in Gridworld and AnimalAI, DHTM outperforms LSTM and CSCG in online Successor Features formation for navigation tasks.
+Modelling sequential data is one of the essential tasks in artificial intelligence as it has many applications, including decision-making and learning world models [1], natural language processing [2], conversational AI [3], time series analysis [4], and video and music generation [5].
+One of the classical approaches to modelling sequential data is to form a representation that stores and condenses the most relevant information about a sequence, and to find a general transformation rule of this information through the dimension of time [6, 7, 8].
+We refer to the class of algorithms that use this approach as Temporal Memory (TM) algorithms because they essentially model the cognitive ability of complex living organisms to remember the past experience and make future predictions based on that memory [9, 10, 11, 12].
+This paper addresses the problem of online sequence learning for planning and decision-making under uncertainty, which can be formalized as reinforcement learning (RL) for a partially observable Markov decision process (POMDP) [13, 14].
+Inferring the hidden state in a partially observable environment is a sequence modelling problem, since it requires processing a sequence of observations to obtain enough information about the hidden state.
+One of the most efficient representations of hidden states for discrete POMDPs is the Successor Representation (SR), which disentangles hidden states and goals given by the reward function [15, 16, 17].
+The Successor Features framework is an extension of SR to continuous POMDP, using the same idea of value function decomposition, but instead for features of a hidden state [18, 19].
+Temporal memory (TM) algorithms can be used to make cumulative predictions about future states and their features, forming SR or SF.
+This work shows that the proposed algorithm, Distributed Hebbian Temporal Memory (DHTM), can dynamically form SFs for navigation tasks in Gridworld and AnimalAI [20] environments.
+The most prominent TM algorithms, such as a Recurrent Neural Network (RNN) [21] or LSTM [9], use backpropagation to capture data relationships, which is known to be unstable due to recurrent nonlinear derivatives.
+They also require complete data sequences to be available during training.
+Although the gradient vanishing problem can be partially circumvented in a way that Receptance Weighted Key Value (RWKV) [22] or Linear Recurrent Unit (LRU) [23] models do, the problem of online learning, i.e.
+learning on non-stationary data sets, is still a viable issue.
+Unlike graphical models such as HMM [24], RNN models and their descendants also lack a probabilistic theoretical foundation, which is advantageous for modelling sequences captured from stochastic environments [25, 26].
+There is little research on TM models that can be used in fully online adaptive systems interacting with partially observable stochastic environments with access to only one sequence data point at a time, a common case in reinforcement learning [27].
+We propose a Distributed Hebbian Temporal Memory (DHTM) algorithm based on factor graph formalism and a multi-compartment neuron model inspired by the HTM neuron [28].
+The resulting graphical structure of our model is similar to that of the factorial HMM [29].
+An important feature of our model is that the transition matrices for each factor are stored as different components (segments) of artificial neurons, which makes computation very efficient in the case of sparse transition matrices.
+Our TM forms sequence representations entirely online and uses only local Hebbian-like learning rules [30, 31, 32], which avoid the drawbacks of gradient methods and make the learning process much faster than gradient methods.
+This comes at the expense of hidden state generalization, since DHTM’s hidden state uniquely encodes each observation sequence.
+Different trajectories leading to the same position in the partially observable Gridworld can result in different hidden states inferred by DHTM.
+In essence, this makes DHTM a model of episodic memory [33], or an adaptive trajectory buffer, with built-in machinery for making multistep future predictions.
+In this light, we also show that an agent forming SF with DHTM can be interpreted as a kind of episodic control [34, 35, 36, 37].
+The proposed TM is tested as a world model for an RL agent architecture navigating in a Gridworld environment and a more challenging AnimalAI testbed [38].
+The results demonstrate that our algorithm outperforms classical LSTM and a biologically inspired HMM-like world model CSCG [39] in changing environments and is in order of magnitude more sample efficient than Dreamer V3 [40].
+Another advantage of our algorithm is that it can be implemented on neuromorphic processors using only local learning rules.
+Our contribution to this work includes the following key elements:
+  • A novel episodic memory model, called DHTM, which is based on a factor graph formalism and features an efficient, bio-inspired implementation.
+  • We demonstrate that the DHTM can operate fully online, using only local Hebbian-like learning rules.
+  • We show that DHTM learns at a speed comparable to model-based table episodic control while supporting distributed representations. This capability significantly broadens its applicability.
 2 Background
-This section provides basic information about some concepts necessary to follow the paper.
+This section provides basic information about some of the concepts necessary to follow the paper.
 2.1 Reinforcement Learning
-This paper considers decision-making in a partially observable environment, usually formalized as a Partially Observable Decision Process (Poupart, 2005).
+This paper considers decision-making in a partially observable environment, usually formalized as a partially observable decision process [13].
 A POMDP is defined as a tuple $\mathcal{M}=(S,A,P,R,O,D,\gamma)$, where $S$—state space, $A$—action space, $P(s,a,s`)=Pr(s`\ |\ s,a)$—transition function, $R(s)$–reward function, O—observation space, $D(a,s`,o)=Pr(o\ |\ a,s`)$—sensor model and $\gamma\in[1,0)$—discount factor, given a transition $s,a\to s`$, where $s\in S$, $a\in A$, $o\in O$.
-If $S,A,O$ are finite, $P,D$ can be viewed as real valued matrices, otherwise, they are conditional density functions.
-Here we consider deterministic rewards, which depend only on the current state, i.e.
+If $S,A,O$ are finite, $P,D$ can be viewed as real-valued matrices, otherwise, they are conditional density functions.
+Here we consider deterministic rewards that depend only on the current state, i.e.
 $R(s):S\to\mathbb{R}$.
 […]
+One of the drawbacks of value-based methods is that they require learning a new Q-function when the reward function changes, even if the environment remains the same.
+To get around this, the dynamics of the environment and the reward function can be decoupled, which leads us to the successor representation framework.
+2.2 Successor Representation
+Successor representations are such representations of hidden states from which we can linearly infer the state value given the reward function [15].
+We assume that the observation and state spaces are discrete.
+$V(h_{t}=i)=\mathrm{E}[\sum^{\infty}_{l=0}\gamma^{l}R_{t+l+1}\ |\ h_{t}=i]=\sum_{j}\mathrm{SR}_{ij}R_{j},$ (1)
+where $\gamma$ is a discount factor, $i$-th row in the matrix $\mathrm{SR}$ is the Successor Representation of a state $i$, and $\mathrm{SR}_{ij}=\sum^{\infty}_{l=0}\gamma^{l}p(h_{t+l+1}=j\ |\ h_{t}=i)$.
+$R_{j}$ is a reward for observing the state $j$.
+An extension of SR for large or continuous state spaces that we use is Successor Features [18], which is based on a similar idea of value function decomposition, but in a feature space instead (see details in Appendix A).
+It is easy to show that, similarly to the Q-function, both SR and SF can be obtained using dynamic programming.
+However, this limits its use to relatively stable environments, since changes in the transition function can’t be taken into account immediately.
+The next step to improve an agent’s adaptability is to use a world model to approximate the environment’s transition function and use it directly to form SR according to its definition.
+And one of the most basic such world models is the Hidden Markov Model.
 2.3 Hidden Markov Model
-Partially observable Markov process can be approximated by a Hidden Markov model (HMM) with hidden state space $H$ and observation space $O$.
-$O$ is the same as in $\mathcal{M}$, but $H$ generally is not equal $S$.
-Variables $H_{t}$ represent an unobservable (hidden) approximated state of the environment which evolves over time, and observable variables $O_{t}$ represent observations that depend on the same time step state $H_{t}$, and $h_{t},o_{t}$ are corresponding values of this random variables.
-For the sake of simplicity, we suppose that actions are fully observable and information about them is included into $H_{t}$ variables.
+A partially observable Markov process can be approximated by a Hidden Markov Model (HMM) with hidden state space $H$ and observation space $O$.
+$O$ is the same as in $\mathcal{M}$, but $H$ is generally not equal to $S$.
+Variables $H_{t}$ represent an unobservable (hidden) approximate state of the environment that evolves over time, and observable variables $O_{t}$ represent observations that depend on the same time-step state $H_{t}$, and $h_{t},o_{t}$ are corresponding values of these random variables.
+For simplicity, we assume that actions are fully observable and that information about them is contained in $H_{t}$ variables.
 For the process of length $T$ with state values $h_{1:T}=(h_{1},\ldots,h_{T})$ and $o_{1:T}=(o_{1},\ldots,o_{T})$, the Markov property yields the following factorization of the generative model:
 $p(o_{1:T},h_{1:T})=p(h_{1})\prod_{t=2}^{T}p(h_{t}|h_{t-1})\prod_{t=1}^{T}p(o_{t}|h_{t}).$ (2)
-In case of discrete hidden state, a time-independent stochastic transition matrix can be learned with Baum–Welch algorithm (Baum et al., 1970), a variant of Expectation Maximization algorithm.
-To compute the statistics for the expectation step, it employs the forward-backward algorithm, which is a special case of sum-product algorithm (Kschischang et al., 2001).
-2.3 Successor Representation
-Successor Representations are such representations of hidden states from which we can linearly infer the state value given the reward function (Dayan, 1993).
-Here, we assume observation and state spaces are discrete.
-$V(h_{t}=i)=\mathrm{E}[\sum^{\infty}_{l=0}\gamma^{l}R_{t+l+1}\ |\ h_{t}=i]=\\ =\sum^{\infty}_{l=0}\gamma^{l}\sum_{j}p(h_{t+l+1}=j\ |\ h_{t}=i)R_{j}=\\ =\sum_{j}\sum^{\infty}_{l=0}\gamma^{l}p(h_{t+l+1}=j\ |\ h_{t}=i)R_{j}=\sum_{j}M_{ij}R_{j},$ (2)
-where $\gamma$ is a discount factor, vector $\mathrm{SR}(h=i)=\{M_{ij}\}_{j}$ is a Successor Representation of a state $i$, and $M_{ij}=\sum^{\infty}_{l=0}\gamma^{l}p(h_{t+l+1}=j\ |\ h_{t}=i)$.
-$R_{j}$ is a reward for observing the state $j$.
-That is, SR can be computed by a TM that is able to predict future states.
-TM algorithms effectively predict observations only for a finite time horizon $T$.
-Therefore, in order to learn SR, a technique similar to TD learning in standard RL may be employed:
-$\delta_{ij}$ $=\sum^{T}_{l=0}\gamma^{l}p(h_{t+l+1}=j\ |\ h_{t}=i))+$ (3)
-$+\gamma^{T+1}\sum_{k}M_{kj}p(h_{t+T+1}=k\ |\ h_{t}=i)-M_{ij},$
-$M_{ij}$ $\leftarrow M_{ij}+\alpha\delta_{ij},$ (4)
-where $\alpha\in(0,1)$ is a learning rate, $\delta_{ij}$—TD error for SR.
-In partially observable environments, however, exact state values are not known, therefore we operate with state distributions or so-called belief states (Poupart, 2005), which are inferred from observations.
-In that case, state value and SR are functions of hidden state variable distribution (see details in Appendix B).
-2.4 Sparse Distributed Representations
-In our work, we design our model to operate with sparse distributed representations (SDRs) to reflect the spatiotemporal property of cortical network activity (Perin et al., 2011).
-In the discrete time case, SDR is a sparse binary vector in a high-dimensional space.
-To encode observed dense binary patterns to SDRs, we use a biologically plausible k-WTA (k-winners take all) neural network algorithm called spatial pooler with a Hebbian-like unsupervised learning method (see details in Appendix A).
+In the case of a discrete hidden state variable, a time-independent stochastic transition matrix can be learned with the Baum-Welch algorithm [41], a variant of the expectation maximization (EM) algorithm.
+To compute the statistics for the expectation step, it uses the forward-backward algorithm, which is a special case of the sum-product algorithm [42].
+Since the transition matrix grows quadratically with the size of the hidden state space, it may become infeasible to train the HMM in large state spaces.
+Since it is possible that not all hidden states are dependent on each other, the Factorial HMM can mitigate this problem by partitioning the hidden state spaces into smaller independent subspaces, providing an efficient way to handle distributed representations.
+We base DHTM on ideas from the Baum-Welch algorithm and the Factorial HMM, while adapting them for online learning.
+2.4 Episodic Control
+Episodic Control (EC) is a general computational approach to storing and reusing an agent’s experience, inspired by the notion of episodic memory [33] from psychology.
+[…]
+In this paper, we refer to EC to better explain the principle of the proposed algorithm for solving RL tasks.
 3 Distributed Hebbian Temporal Memory
 3.1 Factor Graph Model
-Distributed Hebbian Temporal Memory is based on the sum-product belief propagation algorithm in a factor graph (see Figure 1).
-Analogously to Factorial-HMM (Ghahramani and Jordan, 1997), we divide the hidden space $H$ into subspaces $H^{k}$.
-There are four sets of random variables (RV) in the model: $H^{i}_{t-1}$—latent variables representing hidden states from the previous time step (context), $H^{k}_{t}$—latent variables for the current time step, $\Phi^{k}_{t}$—feature variables, and $O^{lm}_{t}$—observable variables.
-Except for $O^{lm}_{t}$, all random variables have a categorical distribution.
-In contrast, $O^{lm}_{t}$, are Bernoulli variables because they represent pixels from a binary input image observation.
-RV state values are denoted as corresponding lowercase letters: $h^{i}_{t-1}$, $h^{k}_{t}$, $\varphi^{k}_{t}$, $o^{lm}_{t}$.
+Distributed Hebbian Temporal Memory is based on the sum-product belief propagation algorithm in a factor graph (see Figure 1A).
+Analogous to Factorial HMM [44], we divide the hidden space $H$ into subspaces $H^{k}$.
+There are four sets of random variables (RV) in the model: $H^{i}_{t-1}$—latent variables representing hidden states from the previous time step (context), $H^{k}_{t}$—latent variables for the current time step, and $\Phi^{k}_{t}$–feature variables.
+All random variables have categorical distributions.
+RV state values are denoted by corresponding lowercase letters: $h^{i}_{t-1}$, $h^{k}_{t}$, $\varphi^{k}_{t}$, $o^{lm}_{t}$.
 Each variable $\Phi^{k}_{t}$ is considered independent and has a separate graphical model for increased computational efficiency.
-However, hidden variables of the same time step are statistically interdependent in practice.
-We introduce their interdependence through a segment computation trick that goes beyond the standard sum-product algorithm (see Eq.
-8).
-The model also has three types of factors: $M^{i}_{t-1}$—messages from previous time steps, $F^{k}_{c}$—context factor (generalized transition matrix), $F^{k}_{e}$—emission factor.
-We assume that messages $M^{i}_{t-1}$ include posterior information from the time step $t-1$, therefore we don’t depict observable variables for previous time steps in Figure 1.
-Figure 1: Partial factor graph for the DHTM.
-The input to the model is a sequence of binary images, each pixel is modelled as Bernoulli random variable $O^{lm}_{t}$, where $l$ and $m$ denote corresponding rows and cols of the image.
-The encoder block forms image categorical features $\Phi^{k}_{t}$ in an unsupervised manner.
+However, in practice, hidden variables of the same time step are statistically interdependent.
+We introduce their interdependence through a segment computation trick that goes beyond the standard sum-product algorithm (see equation 6).
+The model also has three types of factors: $M^{i}_{t-1}$—messages from previous time steps, $F^{k}_{c}$—context factor (generalized transition matrix), $F^{k}_{e}$–emission factor.
+We assume that messages $M^{i}_{t-1}$ contain posterior information from time step $t-1$, so we don’t plot observable feature variables for previous time steps in Figure 1A.
+Figure 1: A.
+Example of a factor graph for the DHTM.
+The input to the model is a sequence of observations $o_{t}$.
+The encoder block forms categorical features $\Phi^{k}_{t}$.
 Each feature $\Phi$ has its own explaining hidden variable, which may depend on hidden variables of the other features and on itself from the previous time step.
@@
 Unary factors $M^{i}_{t-1}$ called messages represent accumulated information about previous time steps.
-Further, we discuss only the upper block of the graph in Figure 1, which is DHTM itself.
-The lower block—an encoder—is described in the Appendix A.
+B.
+Neural implementation.
+Random variables are represented by cell clusters (white circles), where each cell corresponds to a state and its spike frequency to the probability of the state $p(h^{k}_{t})$.
+The dendritic segments of the cell correspond to the context factor values $f_{l}$ for a given combination of states (active presynaptic cells).
+Segments’ excitations $E_{l}$ are combined to determine the cell’s spike frequency $p(h^{k}_{t})$.
+Segment synaptic weights reflect the specificity of the combination of presynaptic cells for the segment.
+Emission factors $F^{k}_{e}$ are fixed and represented by minicolumns within a variable.
+We will discuss only the upper block of the graph in Figure 1A, which is DHTM itself.
+The lower block—an encoder—is described in Appendix C.
 The only requirement for the encoder is that its output should be represented as states of categorical variables (features) for the current observation.
-3.2 Neural Implementation
-The main routine of the DHTM is to estimate distributions of currently hidden state variables given by the equation 5, the computational flow of which is schematically depicted in Figure 2:
+The main routine of the DHTM is to estimate distributions of hidden state variables given by the equation 3, the computational flow of which is schematically shown in Figure 1B:
 $p(h_{t}^{k})\propto\sum_{\Omega_{k}}\prod_{i\in\omega_{k}}M_{t-1}^{i}(h_{t-1}^{i})F_{c}^{k}(h_{t}^{k},\Omega_{k}),$ (3)
 where $\Omega_{k}=\{h^{i}_{t-1}:i\in\omega_{k}\}$, $\omega_{k}={i_{1},\dots,i_{n}}$—set of previous time step RV indexes included in $F_{c}^{k}$ factor, $(n+1)$—factor size.
-Figure 2: Neuronal implementation of the DHTM.
-Random variables are represented by cell clusters (white circles), where each cell corresponds to a state and its spike frequency—to the probability of the state $p(h^{k}_{t})$.
-Cell’s dendritic segments $seg(k)$ correspond to context factor values $f_{l}$ for a particular combination of states (active presynaptic cells) $rec(l)$.
-Segments’ excitations $E_{l}$ are combined to determine cell’s spike frequency $p(h^{k}_{t})$.
-Segment’s synaptic weights reflect specificity of $rec(l)$ combination for the segment.
-Emission factors $F^{k}_{e}$ are fixed and represented by minicolumns inside a variable.
-For computational purposes, we translate the problem to the neural network architecture with Hebbian-like learning (for biological interpretation of the model, see Appendix C).
-As can be seen from Figure 2, every RV can be viewed as a set of spiking neurons representing the RV’s states, that is, $p(h^{k}_{t})=p(c^{j}_{t}=1)$, where $j$—index of a neuron corresponding to the state $h^{k}_{t}$.
-Cell activity is binary $c^{j}_{t}\in\{0,1\}$ (spike/no-spike), and the probability might be interpreted as a spike rate.
+3.2 Neural Implementation
+For computational purposes, we translate the factor graph to the Hebbian learning neural network architecture (for a biological interpretation of the model, see Appendix B).
+As shown in Figure 1B, each RV can be viewed as a set of spiking neurons representing the states of the RV, i.e.
+$p(h^{k}_{t})=p(c^{j}_{t}=1)$, where $j$–is the index of a neuron corresponding to the state $h^{k}_{t}$.
+Cell activity is binary $c^{j}_{t}\in\{0,1\}$ (spike/no spike), and the probability can be interpreted as a spike rate.
 Factors $F^{k}_{c}$ and $M^{i}_{t-1}$ can be represented as vectors, where elements are factor values for all possible combinations of RV states included in the factor.
-Let’s denote elements of the vectors as $f_{l}$ and $m_{u}$ correspondingly, where $l$ corresponds to a particular combination of $k,h^{k}_{t},h^{i_{1}}_{t-1},\dots,h^{i_{n_{l}}}_{t-1}$ state values and $u$ indexes all neurons representing states of previous time step RVs.
-Drawing inspiration from biological neural networks, we group a neuron’s connections into dendritic segments.
-A segment acts as an independent computational unit that detects a particular input pattern (a context state) defined by its own receptive field.
-In our model, a segment links together factor value $f_{l}$, the computational graph shown in Figure 2, and the excitation $E_{l}$ induced by the segment $l$ to the cell it is attached to.
-The segment is active, i.e., $s_{l}=1$ if all its presynaptic cells are active; otherwise, $s_{l}=0$.
-Computationally, a segment transmits its factor value $f_{l}$ to a cell it is attached to if the context matches the corresponding state combination.
+Let’s denote the elements of the vectors as $f_{l}$ and $m_{u}$ respectively, where $l$ corresponds to a particular combination of $k$ state values $h^{k}_{t},h^{i_{1}}_{t-1},\dots,h^{i_{n_{l}}}_{t-1}$, and $u$ indexes all neurons representing states of RVs from previous time steps.
+Inspired by biological neural networks, we group the connections of a neuron into dendritic segments.
+A segment acts as an independent computational unit that detects a particular input pattern (a contextual state) defined by its receptive field.
+In our model, a segment connects the factor value $f_{l}$ and the excitation $E_{l}$ induced by the segment $l$ to the cell to which it is connected.
+The segment is active, i.e.
+$s_{l}=1$ if all its presynaptic cells are active, otherwise $s_{l}=0$.
+Computationally, a segment transmits its factor value $f_{l}$ to the cell to which it is attached if the context matches the corresponding state combination.
 We can now rewrite equation 3 as the following:
@@
 where $L_{l}=\prod_{u\in\mathrm{rec}(l)}m_{u}$ is segment’s likelihood as long as messages are normalized to probability distributions, $\mathrm{seg}(j)$—indexes of segments that are attached to cell $j$, $\mathrm{rec}(l)$—indexes of presynaptic cells that constitute receptive field of a segment with index $l$.
-Initially, all factor entries are zero, meaning cells have no segments.
-As learning proceeds, new non-zero connections grouped into segments are grown.
-In equation 6 we benefit from having sparse factor value vectors because its complexity depends linearly on the amount of non-zero components.
-And that’s usually the case in our model due to one-step Monte-Carlo learning and specific form of emission factors $F^{k}_{e}$:
-$F^{k}_{e}(h^{k}_{t},o^{k}_{t})=\mathbb{I}[h^{k}_{t}\in\mathrm{col}(\varphi^{k}_{t})],$ (7)
-where $\mathbb{I}$—indicator function, $\mathrm{col}(\varphi^{k}_{t})$ is a set of hidden states connected to the feature state $\varphi^{k}_{t}$ that forms a column.
-The form of emission factor is inspired by presumably columnar structure of the neocortex and was shown to induce sparse transition matrix in HMM (George et al., 2021).
-Segment likelihood $L_{l}$, resulting from the sum-product algorithm, is calculated as if presynaptic cells are independent.
-However, it’s not usually the case for sparse factors.
-To take into account, approximately, their interdependence, we substitute the following equation for segment log-likelihood:
-$\log L_{l}$ $=w_{l}\log\frac{1}{n_{l}}\sum_{u\in\mathrm{rec}(l)}w_{ul}m_{u}+$
-$+\sum_{u\in\mathrm{rec}(l)}(1-w_{ul})\log m_{u},$ (8)
-where $w_{pl}$—synapse efficacy or neuron specificity for segment, such that $w_{ul}=p(s_{l}=1|c^{u}_{t-1}=1)$, $w_{l}=\frac{1}{n_{l}}\sum_{u}w_{ul}$—average synapse efficacy for segment $l$, and $n_{l}$-number of cells in segment’s receptive field.
-The idea that underlies the formula is to approximate between two extreme cases:
-  • $p(s_{l}=1|c^{u}_{t-1}=1)\to 1$ for all $u$, which means that all cells in the receptive field are dependent and are part of one cluster, i.e., they fire together.
-In that case, it should be $p(s_{l})=m_{u}$ for any $u$, but we also reduce prediction variance by averaging between different $u$.
-  • $p(s_{l}=1|c^{u}_{t-1}=1)\to 0$ for all $u$ means that presynaptic cells don’t form a cluster.
-In that case, segment activation probability is just a product of the activation probability of each cell.
+Initially, cells have no segments.
+As learning progresses, new non-zero connections, grouped into segments, are created.
+In equation 4, we benefit from having sparse factor value vectors because their complexity depends linearly on the number of non-zero components.
+And that’s usually the case in our model due to the one-step Monte Carlo learning and the special form of the emission factors $F^{k}_{e}$:
+$F^{k}_{e}(h^{k}_{t},\varphi^{k}_{t})=\mathbb{I}[h^{k}_{t}\in\mathrm{col}(\varphi^{k}_{t})],$ (5)
+where $\mathbb{I}$—indicator function, $\mathrm{col}(\varphi^{k}_{t})$ is a set of hidden states connected to the feature state $\varphi^{k}_{t}$ forming a column.
+The shape of the emission factor is inspired by the presumed columnar structure of the neocortex and has been shown to induce a sparse transition matrix in HMM [39].
+The segment likelihood $L_{l}$ resulting from the sum-product algorithm is computed as if the presynaptic cells were independent.
+However, this isn’t usually the case for sparse factors.
+To account for their interdependence, we substitute the following equation for segment log-likelihood:
+$\log L_{l}$ $=w_{l}\log\frac{1}{n_{l}}\sum_{u\in\mathrm{rec}(l)}w_{ul}m_{u}+\sum_{u\in\mathrm{rec}(l)}(1-w_{ul})\log m_{u},$ (6)
+where $w_{ul}$—synapse efficacy or neuron specificity for segment, such that ${w_{ul}=p(s_{l}=1|c^{u}_{t-1}=1)}$, $w_{l}=\frac{1}{n_{l}}\sum_{u}w_{ul}$—average synapse efficacy for segment $l$, and $n_{l}$-number of cells in segment’s receptive field.
+The idea behind the formula is to approximate between two extreme cases.
+The first is ${p(s_{l}=1|c^{u}_{t-1}=1)}\to 1$ for all $u$, which means that all cells in the receptive field are dependent and part of a cluster, i.e.
+they fire together.
+In this case, ${p(s_{l})=m_{u}}$ for all $u$, but we also reduce the prediction variance by averaging across different $u$.
+In the opposite case, ${p(s_{l}=1|c^{u}_{t-1}=1)\to 0}$ for all $u$, presynaptic cells don’t cluster and the segment activation probability is just a product of the activation probability of each cell.
 The resulting equation for belief propagation in DHTM is the following:
 $p(h^{k}_{t})=p(c^{j}_{t}=1)=\underset{j\in\mathrm{cells}[H^{k}_{t}]}{\mathrm{softmax}}(\underset{l\in seg(j)}{\mathrm{max}}(E_{l})),$ (7)
-where $E_{l}=\log f_{l}+\log L_{l}$, $\mathrm{cells}[H^{k}_{t}]$—indexes of cells that represent states for $H^{k}_{t}$ variable.
-Here, we also approximate logarithmic sum with $\mathrm{max}$ operation inspired by the neurophysiological model of segment aggregation by cell (Stuart and Spruston, 2015).
-The next step after computing $p(h^{k}_{t})$ distribution parameters is to incorporate information about current observations $p(h^{k}_{t}\ |\ o^{k}_{t})\propto p(h^{k}_{t})\mathbb{I}[h^{k}_{t}\in\mathrm{col}(o^{k}_{t})]$.
-After that, the learning step is performed.
-The step for closing the loop of our TM algorithm is to assign the posterior for the current step $p(h^{k}_{t}\ |\ o^{k}_{t})$ to $M^{i}_{t-1}$.
+where ${E_{l}=\log f_{l}+\log L_{l}}$, $\mathrm{cells}[H^{k}_{t}]$—indexes of cells representing states for the variable $H^{k}_{t}$.
+Here we also approximate the logarithmic sum with the $\mathrm{max}$ operation, inspired by the neurophysiological model of segment aggregation by cells [45].
+The next step after computing the $p(h^{k}_{t})$ distribution parameters is to incorporate information about the current feature states ${p(h^{k}_{t}\ |\ \varphi^{k}_{t})\propto p(h^{k}_{t})\mathbb{I}[h^{k}_{t}\in\mathrm{col}(\varphi^{k}_{t})]}$.
+After that, the learning step is executed.
+The loop closing step of our TM algorithm is to assign the posterior of the current step ${p(h^{k}_{t}\ |\ \varphi^{k}_{t})}$ to $M^{i}_{t-1}$.
+3.3 Learning
 DHTM learns $f_{l}$ and $w_{ul}$ weights by Monte-Carlo Hebbian-like updates.
-First, $h^{i}_{t-1}$ and $h^{k}_{t}$ are sampled from their posterior distributions: $p(h^{i}_{t-1}\ |\ o^{i}_{t-1})\propto M^{i}_{t-1}$ and $p(h^{k}_{t}\ |\ o^{k}_{t})$ correspondingly.
-Then $f_{l}$ is updated according to the segment’s $s_{l}$ and its cell’s $c^{j}_{t}$ activity so that $f_{l}$ is proportional to several coincidences $s_{l}=c^{j}_{t}=1$ during the recent past, i.e., cell and its segment are active at the same time step:
+First, $h^{i}_{t-1}$ and $h^{k}_{t}$ are sampled from their posterior distributions: ${p(h^{i}_{t-1}\ |\ \varphi^{i}_{t-1})\propto M^{i}_{t-1}}$ and ${p(h^{k}_{t}\ |\ \varphi^{k}_{t})}$ respectively.
+Then $f_{l}$ is updated according to the activity $s_{l}$ of the segment and the activity $c^{j}_{t}$ of its cell, so that $f_{l}$ is proportional to several coincidences ${s_{l}=c^{j}_{t}=1}$ in the recent past:
 $\Delta f_{l}=\alpha(c^{j}_{t}-f_{l})s_{l},$ (8)
-where $\alpha\in[0,1)$ is segment’s learning rate.
-That is, $f_{l}$ is trained in that way that it is equal to the average cell’s activation frequency, given the segment is active.
-It’s similar to Baum-Welch’s update rule (Baum et al., 1970) for the transition matrix in HMM, which, in effect, counts transitions from one state to another, but, in our case, the previous state (context) is represented by a group of RVs, not just one hidden RV.
+where $\alpha\in[0,1)$ is the segment’s learning rate.
+That is, $f_{l}$ is trained to be equal to the average cell activation frequency when the segment is active.
+It’s similar to Baum-Welch’s update rule [41] for the transition matrix in HMM, which in effect counts transitions from one state to another, but in our case, the previous state (context) is represented by a group of RVs.
 Weights $w_{ul}$ are also updated by the Hebbian rule to reflect the specificity of a presynaptic $u$ for activating a segment $l$.
-That is, they are targeted to represent probability $p(s_{l}=1\ |\ c^{u}_{t-1}=1)$ that segment $s_{l}$ is active, given cell $u$ was active at the previous time-step.
-We could learn it by counting activation coincidences and mismatches.
-But in our algorithm it is approximated as exponential moving average of segment’s $s_{l}$ frequency activation, given $c^{u}_{t-1}=1$: $\Delta w_{ul}=\beta\cdot\mathbb{I}[c^{u}_{t-1}=1]\cdot(\mathbb{I}[s_{l}=1]-w_{ul})$, where $\beta\in[0,1)$ — learning rate.
+That is, they are targeted to represent the probability ${p(s_{l}=1\ |\ c^{u}_{t-1}=1)}$ that segment $s_{l}$ is active, given that cell $u$ was active at the previous time step.
+We could learn this by counting activation matches and mismatches.
+But in our algorithm, it is approximated as an exponential moving average of the frequency activation of segment $s_{l}$, given $c^{u}_{t-1}=1$:
+$\Delta w_{ul}=\beta\cdot\mathbb{I}[c^{u}_{t-1}=1]\cdot(\mathbb{I}[s_{l}=1]-w_{ul}),$ (9)
+where $\beta\in[0,1)$ — learning rate.
 3.4 Agent Architecture
 We incorporate DHTM as a part of an RL agent.
-We use the same agent for other memories tested as well.
+The same agent is used for other memories tested.
 The agent consists of a memory model and a feature reward function.
-The memory model generates SF by predicting cumulative future distributions of feature variables $\Phi^{k}$ according to equation 12.
-Here, we assume that feature representations of observations are good enough to define the reward function $r(\varphi^{k}_{t})$, which is also learned during interaction with the environment and, combined with SF representations, is used to estimate the action value function.
-For more details on employed value function decomposition, see Appendix B.
+The memory model generates SF by predicting cumulative future distributions of feature variables $\Phi^{k}$.
+We assume that the reward function can be decomposed linearly in feature space as ${R_{t}=\frac{1}{n}\sum^{n}_{k=1}{r^{k}_{i}}}$, where $r^{k}_{i}$ is a reward associated with feature state $\varphi^{k}_{t}=i$ and $n$—number of feature variables.
+$r^{k}_{i}$ is learned during interaction with the environment and is used in combination with SF representations to estimate the action value function (see Appendix A for details).
 Algorithm 1 General agent training procedure
@@
 The agent training procedure is outlined in Algorithm 1.
-For each episode, the memory state is reset to a fixed initial message with RESET_MEMORY() and action variable is set to a fixed initial action.
-An observation image returned by an environment (obs) is first preprocessed to get events, mimicking a simple event-based camera with a floating threshold determined from the average difference between the current and previous step image intensities.
-The resulting events are encoded to SDRs with a biologically inspired spatial pooling encoder described in Appendix A.
-In OBSERVE() routine, the memory learns to predict next feature states as described in Section 3.
-An agent learns associations to feature states and rewards in REINFORCE() function:
+For each episode, the memory state is reset to a fixed initial message with RESET_MEMORY() and the variable action is set to a fixed initial action.
+An observation returned by an environment (obs) is encoded as a set of categorical variables.
+In the OBSERVE() routine, memory learns to predict next feature states as described in section 3.
+An agent learns associations to feature states and rewards in the REINFORCE() function:
 $r^{k}_{i}\leftarrow r^{k}_{i}+\alpha\mathbb{I}[\varphi^{k}_{t}=i](R_{t}-r^{k}_{i})$ (10)
 where $\alpha$ is a learning rate, $R_{t}$—a reward for the current time step.
-We include actions in the model by forcing some of the hidden variables $H^{k}_{t}$ to represent actions.
-That is, we assume that information about action is included in the hidden state of the model.
-For example, if we have 4 actions, we set 4 states for one of the hidden variables and set its state from observation of the action.
-In SAMPLE_ACION(), for each action and current belief state, TM predicts the next belief states for which SFs of horizon $T$ are formed using uniform strategy, i.e.
-setting action variable to uniform distribution.
-In most experiments, $T$ is dynamically determined from $\Phi^{k}_{t+l}$ distribution.
-If the distribution of $\Phi^{k}_{t+l}$ is close to uniform or states with high positive reward have high probability, the prediction cycle is stopped, rendering the prediction process similar to breadth-first search.
-$\mathrm{SF}_{t+T}(\varphi^{k}=j\ |\ o_{0:t},a_{t})=\\ =\sum^{T}_{l=0}\gamma^{l}\sum_{i}p(\varphi^{k}_{t+l+1}=j\ |\ h^{k}_{t}=i)p(h^{k}_{t}=i\ |\ o_{0:t},a_{t})$ (12)
-$Q(o_{0:t},a_{t})=\sum_{jk}\mathrm{SF}_{t+T}(\varphi^{k}=j\ |\ o_{0:t},a_{t})r^{k}_{j},$ (13)
-where $o_{0:t}$—observation history.
+We include actions in DHTM by forcing some of the hidden variables $H^{k}_{t}$ to represent actions.
+That is, we assume that information about action is contained in the hidden state of the model.
+For example, if we have 4 actions, we set 4 states for one of the hidden variables and set its state from observing the action.
+In SAMPLE_ACION(), for each action and current state, TM predicts feature states up to time step $T$ under policy $\pi$, and the predictions are combined to form SF (equation 11).
+For simplicity, we always build SF under a uniform policy, which doesn’t guarantee optimality, but we found that it doesn’t make a significant difference to iterative policy improvement in our experimental setups.
+The planning horizon $T$ is determined dynamically from the distribution of $\Phi^{k}_{t+l}$.
+If the distribution of $\Phi^{k}_{t+l}$ is close to uniform or states with positive reward have high probability, the prediction cycle is stopped.
+The action value is determined by combining the corresponding SF and feature state rewards as shown in equation 12.
+Finally, the action is sampled from the softmax distribution over the action values.
+$\mathrm{SF}^{\pi}_{t+T}(\varphi^{k}=j\ |\ h_{t})=\sum^{T}_{l=0}\gamma^{l}p_{\pi}(\varphi^{k}_{t+l+1}=j\ |\ h_{t})$ (11)
+$Q^{\pi}(h_{t},a_{t})=\sum_{jk}\sum_{h_{t+1}}\mathrm{SF}^{\pi}_{t+T}(\varphi^{k}=j\ |\ h_{t+1})\cdot p(h_{t+1}\ |\ h_{t},a_{t})\cdot r^{k}_{j},$ (12)
+3.5 DHTM and Episodic Control
+Algorithm 2 Episodic memory learning
+0:  $\varphi_{t+1}$, $a_{t}$
+1:  $h_{t+1}$leftarrow $\mathcal{D}_{a_{t}}(h_{t})$
+[…]
+3:  if $h_{t+1}$ is null or $\varphi^{*}_{t+1}$ is not $\varphi_{t+1}$ then
+4:     $h_{t+1}$leftarrow $\underset{h\not\in K}{\mathrm{argmax}\ }F_{e}(h,\varphi_{t+1})$ $K$ is the set of keys for all actions
+5:     $\mathcal{D}_{a_{t}}(h_{t})$leftarrow $h_{t+1}$
+6:  end if
+7:  $h_{t}$leftarrow $h_{t+1}$
+It can be shown that DHTM combined with the agent architecture described in the previous section is similar to episodic control.
+DTHM performs best when its hidden state space is large.
+Since the hidden state for unpredictable observations is formed randomly, it’s important that there are no accidental state collisions when remembering transitions.
+That is, DHTM tries to remember new sequences without interfering with old memories.
+In effect, this makes DHTM a buffer of agent trajectories with a built-in recall engine.
+Episodic Control is based on a similar idea, and to investigate its difference from DHTM, we implemented an Episodic Control agent, a simplified counterpart of DHTM, reflecting its core property that presumably helps it solve RL tasks.
+In our EC version, $K_{a}$ is an array of hidden states $h_{t}$ and $V_{a}$ is an array of corresponding subsequent hidden states $h_{t+1}$.
+If the key state $h_{t}$ already exists in the dictionary, its value is rewritten.
+That is, $\mathcal{D}_{a}$ represents the transition matrix of a deterministic HMM for a feature variable $\Phi_{t}$.
+The hidden state $h_{t}$ and the feature state $\varphi_{t}$ have the same dependence as in DHTM (equation 5), but here we assume infinite cells per column.
+To predict the next state, we just have to look up $\mathcal{D}_{a}$ for the current state $h_{t}$, but if there is no entry for $h_{t}$ or the prediction does not match the observed state $\varphi_{t+1}$, then the new state is formed.
+Unlike DHTM, the new state is not completely random, but to avoid collisions, the algorithm makes sure that this state hasn’t been chosen before.
+The learning procedure of the EC agent’s memory $\mathcal{D}_{a}$ is summarized in the algorithm 2.
+Since $\mathcal{D}_{a}$ can be transformed into a transition matrix, the process of forming SF can be reduced to one in Sec.
+3.4.
+However, since the resulting transition matrix is sparse, it’s more efficient to consider $\mathcal{D}_{a}$ as a graph representation and to perform SF formation as a breadth-first search (see Appendix D.3 for details).
 4 Experiments
-We test our model in a reinforcement learning task in Gridworld environment, where successor features are easy to interpret and the encoder block (see Sec.
-3) is not required, and in a more challenging AnimalAI 3D environment.
-This section shows how different memory models affect SF learning and an RL agent’s adaptability.
-In our work, we compare the proposed DHTM model with LSTM (Hochreiter and Schmidhuber, 1997) and CSCG (George et al., 2021).
-For each setup, the results are averaged over at least five experiments with different seed values and standard deviation is shown.
-See Appendix D for details about baselines, their model parameters and training regimes.
-In Appendix E, environments and setups are more thoroughly described.
-The source code of experiments and algorithms’ implementations is available at https://anonymous.4open.science/r/dhtm-D02F.
+We test our model in a reinforcement learning task in the Gridworld environment and in a more challenging AnimalAI 3D environment.
+This section shows how different memory models affect the adaptability of an RL agent.
+For each setup, the results are averaged over five experiments with different seed values, and the standard deviation is shown.
+See Appendix D for details on the baselines, their model parameters, and training regimes.
+See Appendix E for a more detailed description of the environments and setups.
+The source code of the experiments and algorithms is at https://anonymous.4open.science/r/dhtm-FA0E.
 4.1 Gridworld
-First we test DHTM in MDP setup, that is, observation state corresponds to position in 5x5 maze.
-Each trial consists of two phases.
-In the first phase, lasting 500 episodes, there are no obstacles on the way to the terminal state that gives a reward to the agent.
-In the second phase, the terminal state is surrounded by a wall, so that the agent has to detour in order to get the reward, which remains in the same position.
-Each episode starts with the agent in the same position and ends when terminal state is entered (i.e.
-the reward is collected) or maximum action steps reached.
-Each episode, we measure the amount of action steps for agent required to reach the goal.
-The results show (Fig.
-3) that the agent that uses DHTM for SR formation faster converges to the optimal trajectory than CSCG.
-CSCG is able to converge to optimal trajectory only if the initial training dataset is gathered using random strategy after environmental changes.
-It requires an oracle that tells when changes occurred to gather a new training dataset.
-Otherwise, the agent with CSCG memory converges to suboptimal trajectories as the training batch gets very correlated trajectories following the current agent’s policy, which causes overfitting.
-Here, we also provide results of Q-table and SR-table agents for reference.
-Figure 3: Amount of steps required for the agent to reach the goal in fully observable 5x5 Gridworld using different types of memory: DHTM and CSCG.
-CSCG is trained in two modes.
-The first mode is with oracle, which tells the agent to gather some trajectories using uniform policy after environment changes.
-The second mode is without oracle, so the agent continues using its current strategy to gather new trajectories.
-At the 500th episode, the terminal state is blocked by a wall, so the agent needs to find a new way to the goal.
-There are also plots of learning curves for Q-table and SR-table agents for reference.
-The second test was conducted in partially observable Gridworld in a 5x5 changing maze with a stationary terminal state as before.
-However, now all the agent can see is the colour of the cell in which it is located.
-The agent is also able to see the colour of an obstacle it encounters, receiving additional punishment.
-As can be seen from Figure 4, in contrast to DHTM, CSCG and LSTM are not able to converge after reward is blocked at the 300th episode.
-Even with an oracle signalling about changes in the environment, which enables to gather new fully random trajectories, neither CSCG nor LSTM are able to relearn maze’s structure in online case scenario.
-It can be seen, however, that DHTM converges to a suboptimal trajectory of length 13 on average vs 9 steps for optimal trajectory, after reward is blocked.
-This may be the result of inferior generalisation properties of DHTM, which we discuss in Section 5.
-Figure 4: Learning curves measured in action steps until the goal for the agent using different memory types for SF formation in partially observable 5x5 Gridworld environment.
+Figure 2: POMDP 5x5 Gridworld foraging task using different types of memory: DHTM, CSCG, LSTM.
+CSCG and LSTM are trained in two modes.
+The first mode is with oracle, which tells the agent to reinitialize memory and clear the trajectory buffer during environmental change.
+The second mode is without oracle, so the agent continues using its current strategy to gather new trajectories and update memory.
+There is also an episodic control agent (ec) among baselines, which uses a simple dictionary to store trajectories and make predictions.
+A.
+Dynamic of steps required for the agent to reach the goal.
 In the 300th episode, the goal is blocked by a wall, so the agent needs to find a new optimal trajectory.
-CSCG is trained in two modes.
-The first mode is with an oracle, which tells the agent to gather some trajectories using uniform policy after the environment changes.
-The second mode is without an oracle, so the agent continues using its current strategy to gather new trajectories.
-In these experiments, we aimed to show that in cases of quite limited and non-stationary trajectory buffer size, an RL agent armed with DHTM memory is more adaptive than those using LSTM and CSCG.
-However, we do not claim that it is impossible, in principle, to train LSTM and CSCG in such a setup.
-We only show that DHTM notoriously fits this kind of task without numerous workarounds that would be required by those classical methods.
+B.
+Total steps taken during 1000 episodes.
+The first test was conducted in partially observable Gridworld in a 5x5 alternating maze with a stationary end state and the agent’s starting position.
+The agent can observe the colour of the cell it is in and the colour of an obstacle it encounters, while being punished in addition to the constant step penalty.
+Since the observation can be represented directly as a categorical variable with the number of states equal to the number of possible colours, no encoder is needed.
+Each episode starts with the agent in the same position and ends when the agent enters the terminal state (i.e., when the reward is collected) or when the maximum number of action steps is reached.
+In each episode, we measure the number of action steps required for the agent to reach the goal.
+Each trial lasts $1000$ episodes and consists of two phases.
+In the first phase, which lasts $300$ episodes, there are no obstacles on the shortest path to the terminal state.
+In the second phase, the previously optimal path is blocked by a wall, so the agent has to take a detour to get the reward.
+The experiment aims to show the ability of a memory to relearn the structure of the maze in the test time.
+If the agent is able to reach the end state in the second phase, it suggests that the model can forget, or at least suppress, the memory of the old shortest path when it is blocked.
+Otherwise, the agent will continue to hit the wall and fail to reach the terminal state.
+In this experiment, we compare our memory model DHTM with CSCG [39] and LSTM [9] within the same agent architecture that uses memory to form successor features and evaluate actions (see Sec.
+3.4).
+CSCG and LSTM are trained in two modes.
+The first mode is with an oracle that tells the agent to reinitialize memory and clear the trajectory buffer just in time for the environmental change.
+The second mode is without an oracle, so the agent continues to use its current strategy to collect new trajectories and update memory in the second phase.
+There is also a simple episodic control agent (EC) among the baselines, which uses a dictionary to store trajectories and make predictions (see Section 3.5 for details).
+The results in Figure 2 show that CSCG effectively can’t adapt to the environmental change in the online regime, but only if it is retrained from scratch on new trajectories.
+And, apparently, LSTM is only able to learn a predictive model for a small horizon, since it performs significantly worse than other baselines in the second phase, even with oracle enabled.
+In contrast, DHTM and EC are able to adapt online without explicit memory reset.
+Another important observation is that the EC agent performs on par with DHTM, suggesting that its mechanism can explain most of DHTM’s performance in this task.
 4.2 AnimalAI
-AnimalAI is a 3D environment that is able to give an agent rich visual input in the form of RGB images.
-The objective of the experiment described here is to give a hint that DHTM is also able to deal with complex visual input due to its encoder block described in Appendix A.
-The agent’s task is to reach a ball, which is food, that is located in the same room, which is 10 by 10 meters.
-Start agent and food positions are the same between episodes.
-All the agent needs to do is find the shortest path to the goal, as in the Gridworld experiments.
-In each episode, we measure the number of actions (steps) the agent takes to reach the goal.
-The results are shown in Figure 5.
-We can see that after 150 episodes with the maximum number of steps set to 40, the agent stably reaches the goal.
-Although slight instability persists, the agent’s path is near optimal.
-We weren’t able to train CSCG and LSTM with the same encoder block as we used for DHTM, even after increasing their trajectory buffer sizes.
-We think that the problem might be because of the non-stationarity of the encoder’s output, which adapts continuously.
-Figure 5: Amount of action steps the agent uses to reach the goal in a 10x10-meter room in AnimalAI environment using DHTM memory for SF formation.
+Figure 3: Amount of action steps the agent performs to reach the goal in a 10x10-meter room in AnimalAI environment using DHTM, CSCG, LSTM and EC dictionary.
 The agent and the goal have the same initial positions throughout the episodes.
-Each timestep, the agent gets an image of a first-person view of the environment and produces one of three actions: turn left, turn right, or go forward.
+Each time-step, the agent gets an encoded image of a first-person view of the environment and produces one of three actions: turn left, turn right, or go forward.
+A.
+Dynamic of steps required for the agent to reach the goal.
+B.
+Total steps taken during 1000 episodes by the agent with DHTM with K-Means encoder and Categorical VAE.
+For the latter, DHTM’s context factor size is varied (i.e., how many previous time-step hidden variables are considered for prediction).
+AnimalAI is a 3D environment capable of providing rich visual input to an agent in the form of RGB images.
+The goal of the experiment described here is to show that DHTM can handle distributed representations and is scalable to environments with larger state spaces.
+[…]
+We test the agent using two different encoders: K-Means and Categorical VAE [46].
+Both encoders are pre-trained on RGB images collected with a uniform strategy during $2000$ episodes in the same room used for agent testing.
+K-Means represents an image as a single categorical variable with $50$ states, and Categorical VAE is configured for distributed representation with five variables with $50$ states each (see Appendix C for details).
+[…]
+To show that DHTM can’t simply be replaced by multiple parallel EC dictionaries or CSCGs in the case of distributed encoding, we also vary the context factor size of DHTM, that is, how many previous time-step variables are taken into account to predict the current state of a variable.
+The factor size of 1 corresponds to the graph of the Factorial HMM, with parallel HMMs.
+It can be seen from Figure 3B that in our case of categorical VAE, apparently knowing only the previous state of the variable is not enough to reliably predict rewarding features, since the agent doesn’t converge to the shortest path.
+We also tested the variant (factor 1*) where each feature variable is represented by three hidden states, to ensure that the result for factor size $1$ is not due to DHTM’s hidden state collisions.
+DHTM with a factor that takes into account more than one previous time-step hidden variable (excluding the action variable) performs significantly better.
+For diagrams depicting factor graphs used in these experiments, see Appendix D.4.
 5 Conclusion
-This paper introduces a novel probabilistic Factorial-HMM-like algorithm, DHTM, for learning an observation sequence model in non-stationary environments that uses local Hebbian-like learning rules.
-DHTM is scalable to multiple feature variables by employing sparse distributed representations and sparse factor function implementation using segments, which inspire biologically plausible multicomponent neural models.
-In contrast to methods that use Monte-Carlo trajectory sampling for future state probability estimation, our method can perform belief propagation, so each prediction step adds a constant amount of computation, allowing a long planning horizon.
-Experiments show that our memory model can quickly learn and unlearn observation sequences and transition dynamics, even in changing environments.
-In that case, methods that rely on backpropagation, like LSTM and CSCG, are difficult to train and require rigorous trajectory buffer management, a common problem in RL tasks.
-One of the main limitations of the DHTM is that its temporal context is completely random, as it is formed on the fly using only the local information available.
-As a result, the mechanism of context formation allows only limited generalizations between different trajectories.
-However, this is what makes fast learning possible.
-Combining this fast learning with more slow learning that allows generalizations is a promising direction.
-Forming Successor Features combined with a two-level hierarchy of DHTM layers may provide the next step to a more generalized sequence model.
-The first (fast) DHTM layer may be used to form Successor Features both for planning and as current state representation, which is then observed by the second (slow) DHTM layer.
-We hypothesize that because SFs should provide a better generalization than the raw features, the second layer should be able to form a more versatile model of the environment.
-6 Impact Statements
-This paper presents research that aims to advance the field of machine learning. There are numerous potential societal consequences of our work, none of which we believe require special emphasis here.
-Appendix A Encoding and Decoding Observations
-Because our model is designed to work with sparse distributed representations and the testing environments do not provide observations as SDRs by default, an encoding procedure is required.
-For this task, we use a modified version of the Spatial Pooler (SP) [Cui et al., 2017, Mnatzaganian et al., 2017], a distributed noise-tolerant online clustering neural network algorithm that converts input binary patterns into SDRs with fixed sparsity while retaining pairwise similarity [Kuderov et al., 2023].
-The SP algorithm learns a spatial specialization of neurons’ receptive fields using the local Hebbian rule and k-WTA ($k$ winners take all) inhibition [Oster et al., 2009].
-Here we outline the main differences from the “vanilla” version of the SP algorithm described in Cui et al.
-[2017].
-During an agent’s decision-making process pipeline, the SP encoder accepts a current observation $o$ and transforms it to a latent state SDR $z$.
-Observation $o$ should be an array of binary states.
-To transform an RGB image into a binary array, we perform the following processing steps, simulating a simple event-based camera:
-  1. 1.
-    Transform image to greyscale.
-  2. 2.
-    Calculate difference between current and previous greyscale image $d$.
-  3. 3.
-    Determine average difference over the entire image $\tau=\mathrm{mean}(d)$.
-  4. 4.
-    For pixels that have $d$ above $\tau$ set $1$, for others set 00.
-In terms of processing, our SP encoder functions as a standard artificial neural network with a k-WTA binary activation function.:
-$\text{overlaps}_{i}$ $=\beta_{i}W_{i}o$ (14)
-$z_{i}$ $=\mathbb{I}\left[i\in\text{kWTA}(\text{overlaps})\right],$ (15)
-where $o$—a binary observation vector, $W_{i}$—a row-vector representing $i$-th neuron’s connection weights (where non-existing connections have zero weights), $\text{overlaps}_{i}$—a value representing the strength of the input pattern recognition with the neuron $i$  [^1] While the name “overlap” does not exactly reflect its meaning in our SP modification, because it is not a binary overlap between a receptive field and an input pattern, we kept it on purpose to refer to the similar term commonly used for the original SP., $\beta_{i}$—an $i$-th neuron boosting value, $z_{i}$—an $i$-th bit of an output SDR, $\mathbb{I}[\dots]$—an indicator function, kWTA—a $k$-winners-take-all activation function returning $k$ indices of the neurons with the highest overlap.
-One difference between the “vanilla” SP algorithm and ours is that we do not distinguish between potential and active neural connections.
-Because all [existing] connections are active, they all participate in calculating overlaps.
-In the overlaps calculation, non-binary, that is, real-valued weights are used, similar to artificial neural networks, as shown in equation 14.
-Furthermore, each neuron has a fixed capacity to produce neurotransmitters, which it distributes between its synaptic connections.
-This means that we keep all neuron weights normalized and summed to one.
-While it achieves the same Hebbian learning with homeostatic plasticity as the original SP, the exact formula is slightly different:
-$\tilde{W}_{i}$ $=W_{i}+\alpha z_{i}\frac{\text{RF}_{i}\odot o}{\sum_{j}{\text{RF}_{i}\odot o}}$
-$W_{i}$ $\leftarrow\frac{\tilde{W}_{i}}{\sum_{j}{\tilde{W}_{j}}},$ (16)
-where $\tilde{W}_{i}$—a row of new $i$-th neuron weights before normalization, $\alpha$—learning rate, $z_{i}$—a binary value representing the current activity state of the $i$-th neuron, $RF_{i}$—an $i$-th row of the binary connectivity matrix representing an $i$-th neuron receptive field, $\odot$—elementwise product, $o$—a binary observation vector.
-The original SP algorithm has several drawbacks, including encoding instability caused by an innate homeostatic plasticity mechanism known as boosting, which helps neurons specialize and increases overall adaptability but makes memorization tasks more difficult, and slow processing on large inputs such as images, where an encoding overhead becomes noticeable when compared to overall model timings around 1k input size.
-The introduction of the newborn stage, which follows the ideas proposed in Dobric et al.
-[2022], solves an encoding instability problem.
-The newborn stage of a spatial pooler occurs during the early stages of its learning process, when its neurons are expected to specialize.
-The boosting, which is intended to aid in the specialization process, is activated only during the newborn stage and its scale gradually decreases from the configured value to zero.
-Boosting remains turned off during an encoder’s “adulthood”, reducing the possibility of spontaneous re-specialization.
-To reduce processing overhead, we use a much more sparsified connection matrix than in the original SP version.
-We randomly initialize connections with 40-60% sparsity, which is typical for the “vanilla” SP.
-Then, during the newborn stage, we gradually prune the vast majority of the weakest connections, resulting in neurons that are highly specialized due to their small receptive fields.
-We typically configure the final receptive field size in relation to the average input pattern size (usually 25-200% of it, resulting in 0.1-10% connections sparsity).
-For example, if binary input patterns have on average 100 active bits out of 1000, we can set the target size of receptive fields to 25, which is 25% of the active input size and corresponds to 2.5% connection matrix sparsity.
-As a result, the spatial pooler’s instability (and thus adaptiveness!) becomes even more limited in the adult stage.
-Because of its soft discretization (from the distributed representation) and clusterization properties, we expect SP to assist the model with input sequence memorization and an environment transition dynamics generalization tasks in addition to the encoding itself.
-However, because the SP encoder learns online, particularly during the newborn stage, its output representation can be highly unstable during the early stages, potentially resulting in a performance drop.
-To visualize and debug an encoded observation, we also learn a decoder, which is a linear neural layer learned locally with gradient descend on the MSE error between the predicted reconstruction and the actual observation.
-Appendix B Value Function Decomposition
-In our agent model, we approximate the reward function R(s) as a sum: $R_{t}=\frac{1}{n}\sum^{n}_{k=1}{r(\varphi^{k}_{t})\varphi^{k}_{t}}$, where $r(\varphi^{k}_{t})$ is a reward associated with state $\varphi^{k}_{t}$, $n$–number of feature variables.
+This paper introduces a novel episodic memory model, DHTM.
+It combines ideas from belief propagation, episodic control, and neurophysiological models of the neocortex, resulting in a fast probabilistic sequence memory.
+The experiments show that our memory model can quickly learn observation sequences in changing environments and effectively reuse them for RL tasks, on par with model-based table episodic control.
+We also show that, unlike CSCG and table EC, DHTM is able to handle distributed representations and account for statistical dependencies between variables of these representations.
+Potentially, it extends the application of our algorithm to higher state spaces, exploiting their factorizations for more efficient computations and to reduce the number of stored parameters.
+Our results suggest that DHTM can be viewed as an adaptive trajectory buffer for a slower generalizing algorithm, such as a Neural State Space Model-based [47].
+Through prediction error learning, DHTM can selectively discard and replace obsolete transitions in observed trajectories without disrupting relevant experience.
+It can also provide an initial policy for an agent through episodic control until a slower but more accurate world model takes over.
+These features of a trajectory buffer are highly desirable in the case of online learning in non-stationary environments.
+Further improvements of DHTM can be directed towards the implementation of an automatic feature space factorization discovery to determine an optimal context factor size during learning.
+To avoid excessive segment growth in noisy environments, it’s also possible to implement an adaptive segment activation threshold.
+One of the main limitations of this study is that the evaluation of the algorithms is limited to basic RL setups due to the simple agent architecture.
+More advanced use cases could be demonstrated using better exploration strategies and in combination with other memory models, which is also a matter of future research.
+Appendix A Value Function Decomposition
+In our agent model, we approximate the reward function R(s) as a sum:
+$R_{t}=\frac{1}{n}\sum^{n}_{k=1}{r(\varphi^{k}_{t})},$ (13)
+where $r(\varphi^{k}_{t})$ is a reward associated with state $\varphi^{k}_{t}$, $n$–number of feature variables.
 Then, similarly to the Successor Representation idea (see Section 2.2), the value function can be represented as:
 $V(h_{t})$ $=\mathrm{E}[\sum^{\infty}_{l=0}\gamma^{l}R_{t+l+1}\ |\ h_{t}]=\sum^{\infty}_{l=0}\gamma^{l}\mathrm{E}[\frac{1}{n}\sum^{n}_{k=1}{r(\varphi^{k}_{t})}\ |\ h_{t}]$
-$=\frac{1}{n}\sum^{n}_{k=1}\sum_{j}\sum^{\infty}_{l=0}\gamma^{l}p(\varphi^{k}_{t+l+1}=j\ |\ h^{k}_{t})r^{k}_{j}$
-$=\frac{1}{n}\sum^{n}_{k=1}\sum_{j}M^{k}_{j}(h^{k}_{t})r^{k}_{j},$ (17)
-where $M^{k}_{j}(h^{k}_{t})=\sum^{\infty}_{l=0}\gamma^{l}p(\varphi^{k}_{t+l+1}=j\ |\ h^{k}_{t})$, $h_{t}=(h^{1}_{t},...,h^{n}_{t})$—hidden state vector of variables $\{H^{k}_{t}\}_{k}$.
-Appendix C Biological Interpretation
+$=\frac{1}{n}\sum^{n}_{k=1}\sum_{j}\sum^{\infty}_{l=0}\gamma^{l}p(\varphi^{k}_{t+l+1}=j\ |\ h_{t})r^{k}_{j}$
+$=\frac{1}{n}\sum^{n}_{k=1}\sum_{j}\mathrm{SF}^{k}_{j}(h_{t})r^{k}_{j},$ (14)
+where $\mathrm{SF}^{k}_{j}(h_{t})=\sum^{\infty}_{l=0}\gamma^{l}p(\varphi^{k}_{t+l+1}=j\ |\ h_{t})$, $h_{t}=(h^{1}_{t},...,h^{n}_{t})$—hidden state vector of variables $\{H^{k}_{t}\}_{k}$.
+Appendix B Biological Interpretation
 Neural implementation of the DHTM is inspired by neocortical neural networks (see Fig.
-6).
+4).
 Hidden variables $H^{k}$ may be considered as populations of excitatory pyramidal neurons in cortical layer L2/3 of somatosensory areas, with lateral inhibition modelled as $\mathrm{softmax}$ function.
@@
 Messages are propagated through synapses of dendritic segments, which correspond to factors $F^{k}_{c}$.
-Dendritic segments of biological neurons are known to be coincidence detectors of its synaptic input [Stuart and Spruston, 2015].
+Dendritic segments of biological neurons are known to be coincidence detectors of their synaptic input [45].
 We use the notion of dendritic segment to sparsely represent context factors $F^{k}_{c}$, as each factor value corresponds to a particular combination of states (or active cells).
 Feature variables $\Phi^{k}_{t}$ may be considered to represent cells of a granular layer (L4), as they are known to be the main hub for sensory excitation for L2/3.
-L2/3 cells that have common sensory input from the layer L4 are modelled as columns for particular feature states $\mathrm{col(\varphi^{k}_{t})}$ [Mountcastle, 1997].
+L2/3 cells that have common sensory input from the layer L4 are modelled as columns for particular feature states $\mathrm{col(\varphi^{k}_{t})}$ [49].
 Figure 4: Biological view of the neural implementation of the DHTM.
@@
 Spike frequency of a neuron reflects state probability $p(h^{k}_{t})$ of a variable.
+Appendix C Encoding Observations
+Since DHTM is designed to process categorical random variables, we need a categorical encoder for RGB images to solve AnimalAI tasks.
+[…]
+[…]
 Appendix D Implementation details
 As mentioned in Section 3.4, we incorporate DHTM as a part of an RL agent, which has a memory model and a feature reward function.
-All tested memory models share the same pipeline—they learn sequences of encoded binary observations (i.e.
-SDRs that we get from Spatial Pooler encoder described in Appendix A) concatenated with one-hot encoded actions.
+All tested memory models share the same pipeline—conditioned on actions, they learn sequences of categorically encoded observations and are used to form SFs.
+In all experiments, the agent uses softmax exploration with temperature equal to $0.04$, except for EC agent we set softmax temperature to $0.01$, since for higher temperatures its policy is unstable.
+For SF formation the agent uses $\gamma=0.8$ and $\gamma=0.9$ for Gridworld and AnimalAI experiments respectively.
+Maximum SF horizon is set to $50$, but it is usually shorter since we employ early stop condition based on feature variables distribution.
+That is, if the prediction is close to uniform or the probability of a state with associated positive reward meets a threshold, the planning is interrupted.
+We use probability thresholds equal to $0.05$ and $0.01$ for Gridworld and AnimalAI respectively.
 D.1 LSTM
-LSTM baseline was implemented with a single LSTMCell from PyTorch library [Paszke et al., 2019].
+LSTM baseline was implemented with a single LSTMCell from PyTorch library [52].
 It is supported by an additional symexp-layer to encode input before passing it to the LSTM cell and a symexp-layer to decode the LSTM cell’s output from the LSTM’s hidden state back to the input representation, where symexp activation function, $\mathrm{symexp}(x)=\mathrm{sign}(x)e^{\lvert x\rvert-1}$, is a reverse of symlog function: $\mathrm{symlog}=\mathrm{sign}(x)\log{(\lvert x\rvert+1})$.
-For the Gridworld task, LSTM with a hidden state size of 100 is trained on a buffer of 100 trajectories with a decreasing learning rate starting at 0.02.
-25 trajectories are left out for validation on each epoch.
-The training is terminated when the validation loss is not decreasing more than 3 epochs in a row.
-After that, 30% of the buffer is replaced with new trajectories, and training continues.
+For both Gridworld and AnimalAI tasks, LSTM with a hidden state size of 100 is trained on a buffer of $1000$ trajectories with learning rate $0.002$ on $50$ randomly sampled batches of size $50$ every $10$ episodes.
 We also incorporated some notion of random variables in hidden states by splitting the hidden state of the tested LSTM into groups.
 In all experiments the hidden state represents $10$ categorical variables with $10$ states.
-That is, LSTM is forced to learn $10$ categorical distributions with multi-cross-entropy loss to explain the observed sequences, which is a somewhat close to the multi-categorical hidden state representation used in DreamerV2/V3 [Hafner et al., 2023].
+That is, LSTM is forced to learn $10$ categorical distributions with multi-cross-entropy loss to explain the observed sequences, which is a somewhat close to the multi-categorical hidden state representation used in DreamerV2/V3 [40].
 The idea of using symexp activation function, mentioned above, is inspired by Dreamer too, and is used to remedy the problem of learning extreme probability values.
@@
 CSCG baseline was implemented using code from the repository accompanying the paper (https://github.com/vicariousinc/naturecomm_cscg).
-CSCG with 10 clones per observation state was trained on batches with a size of 2500 observation steps for Gridworld tasks with 1000 EM algorithm iterations.
-We iteratively calculated the exponential moving average of transition matrices obtained for different batches with a smoothing coefficient $\alpha=0.9$ as described in the paper [Dedieu et al., 2019].
+CSCG was trained on buffers of $500$ action steps with $1000$ EM steps until convergence.
+We iteratively calculated the exponential moving average of transition matrices obtained for different batches with a smoothing coefficient $\alpha=0.05$ as described in the paper [53].
 This smoothed transition matrix was used as an initialization for the next training batch and for inference.
 The first batch of observations is gathered using the uniform policy, and later batches are gathered using the current best policy.
-In so-called oracle mode, after the environment changes, the agent gathers a uniform policy batch again.
-D.3 DHTM
-In the POMDP Gridworld task, a graph for DHTM consists of three hidden variables $H$ with 40 cells per feature state, which is raw observation with 7 states in the case of Gridworld.
-Each hidden variable gets the same observation, and their predictions are averaged.
+In so-called oracle mode, the trajectory buffer is continuously growing, and the transition matrix is trained from scratch every $500$ steps.
+The buffer is discarded at the environmental change and the transition matrix is reinitialised.
+D.3 Episodic Control
+Here, we provide an algorithm for SF formation using EC dictionary $\mathcal{D}_{a}$ (see Algorithm 3).
+In contrast to other memory algorithms, EC dictionary doesn’t estimate state transition probabilities, but it stores connections between them in all-or-nothing faction.
+Therefore, on each time step, all predicted observations are assumed equally probable.
+But it’s not a limitation in our set of experiments, since we tested our algorithms in deterministic environments.
+Algorithm 3 SF formation for EC
+0:  $h_{0}$, $\gamma\in(0,1)$, $T$
+0:  $\mathrm{SF}$
+1:  $\mathrm{SF}$leftarrow array of zeros
+2:  $\mathrm{PS}$leftarrow {$h_{0}$} # $\mathrm{PS}$ is the set of all states (nodes) on the current BFS depth
+3:  goal_found $\leftarrow$ false
+4:  for $l=1..T$ do
+5:     $\mathrm{PS}$leftarrow $\bigcup_{a\in A}\{\mathcal{D}_{a}(h)\}_{h\in\mathrm{PS}}$ get next depth nodes assuming the policy is uniform
+6:     counts $\leftarrow$ array of zeros
+7:     for all $h\in\mathrm{PS}$ do
+[…]
+9:        countsφ = countsφ + 1
+10:        goal_found $\leftarrow$ $r_{\varphi}>0$ check if the feature state is rewarding
+11:     end for
+12:     $\mathrm{SF}$ = $\mathrm{SF}$ + $\gamma^{l-1}$ NORMALIZE(counts)
+13:     if goal_found is true then
+14:        break
+15:     end if
+16:  end for
+D.4 DHTM
+In the Gridworld task and AnimalAI with K-Means encoder, the factor-graph for DHTM consists of three hidden variables $H$ with $40$ states per one feature state.
+Each hidden variable gets the same copied observation, and their predictions are averaged.
 Using several hidden variables here is essential, as it significantly increases the amount of trajectories memory is able to store by decreasing collision probability for hidden state representations of different trajectories.
-The same graph is used for the AnimalAI task, but with the encoder block forming one feature variable with 50 states.
-In MDP Gridworld, a graph with one hidden variable and one cell for observation state is used, as there is no need to distinguish observations in different contexts.
-Generally, the softmax strategy much better fits DHTM than Epsilon greedy.
-This is connected with the hidden state generalization problem discussed in Section 5.
-Each new trajectory creates a new hidden state, for which possible future trajectories are not known.
-The more exploration there is, the less useful the memory is.
-Softmax renders exploration more narrowed around already-known good trajectories, allowing better reuse of DHTM’s knowledge.
-SFs are formed using $\gamma=0.8$ in all experiments except MDP Gridworld, where $\gamma=0.99$ is used.
-The maximum planning horizon is set to 50 in POMDP Gridworld and AnimalAI, with stopping criteria based on predicted feature variable distribution: 1) DKL between prediction and uniform distribution is less than 0.1, 2) net probability of feature states with positive rewards is greater than 0.01.
-For MDP Gridworld, planning steps are fixed and equal to 10.
-The same SF parameters were used for CSCG and LSTM.
-In all experiments, segments are initialized with the maximum value, which is $1$.
-Synapses are initialized with $0.5$.
-Segments’ and synapses’ learning rates are $0.1$.
-We also prune segments with values less than $0.3$ to speed up learning.
+In AnimalAI experiments, hidden variables state spaces are of size $40$ states per one feature state.
+For Categorical VAE encoder, we use one hidden variable per feature variable, except the case (factor 1*) that tests that information from one feature variable is not sufficient for correct prediction, in which three hidden variables per observation variable are used and the factor of size three connecting to only hidden copies of the same feature variable.
+Examples of typical DHTM’s graphical models used in our experiments are presented on Figure 5.
+It’s important to note that edge colours denote separate graphical models, but the variables are shared (copied) between them.
+So there are no loops in these models.
+For simplicity, we define context factor size as the number of hidden variables from the previous time-step it has connections to, but it also connects to one hidden variable of the current step and the action variable.
+[…]
+Figure 5: Examples of factor-graphs for two main setups.
+Edge colours show that edges correspond to separate graphical models with copied variables.
+All context factors are also connected to action variable $A_{t-1}$.
+A.
+Graphical model for one feature variable.
+In that case, feature value is copied for three hidden variables, to increase hidden state space and avoid memory collisions.
+B.
+Graphical model for Categorical VAE distributed encoding, example for factor of size $3$ (i.e.
+takes into account tree hidden variables from the previous time-step).
 Appendix E Experimental Setups
@@
 There are two special colours for the world border and obstacles.
-Each timestep, an agent gets the colour index of the current agent’s position.
-In MDP mode, the agent gets a tuple of its coordinates in the maze.
-An agent is able to see obstacles or border colour when trying to go through them.
+Each time-step, an agent gets the colour index of the current agent’s position.
+An agent is able to see obstacles or border colour when trying to go through them, which just replace current position colour for one step without actually changing the agent’s position.
 We also punish an agent trying to go through a wall or world border in addition to base step reward punishment to facilitate faster convergence to optimal trajectories.
-Setups used in our experiments are depicted in Figures 7 and 8.
+Samples of setups used in our experiments are depicted in Figure 6.
 Here, the red colour corresponds to obstacles and the light gray colour to the reward.
+Floor colours are resampled for each trial, but we ensure that there are no more than five floor colours, including the terminal state.
 In each episode, the agent and the terminal rewarding state have the same initial positions.
 The episode ends when the terminal state is entered or the limit of action steps is reached.
-Figure 7: Schematic view of MDP Gridworld setups, in which agent gets its Gridworld position instead of colour index.
-The left picture shows the initial setup of the environment, with the reward in the center.
-The right picture represents the environment after the reward is blocked by obstacles (red squares).
-Agent’s initial position in the lower left corner.
 Figure 6: Schematic view of POMDP Gridworld setups.
@@
 E.2 AnimalAI
-AnimalAI is a testbed inspired by experiments with animals [Crosby et al., 2020].
+AnimalAI is a testbed inspired by experiments with animals [38].
 The environment consists of 3D area surrounded by a wall and many different objects that can be placed using a configuration file, including walls, food, ramps, trees, movable obstacles, and more.
 We tested our agent in a 10x10-meter room surrounded by walls (see Fig.
-9).
+7).
 Each timestep, the agent gets an RGB image of a first-person view of the environment and is able to influence its behaviour by choosing one of three actions: turn left, turn right, or go forward.
@@
 Figure 7: Sample of the agent’s first-person view of our setup in AnimalAI environment. The green ball represents a goal, a food item to be collected.
-Appendix F Glossary
+Appendix F Additional Experiments
+In order to ground our experiments to state-of-the-art model based algorithms, as an additional baseline, we test the smallest version of Dreamer V3 with 18 million parameters.
+We use the implementation described in Voudouris et al.
+[54], which is based on Hafner et al.
+[40] and adapted for AnimalAI.
+[…]
+Figure 8: Dreamer V3 (18M) score in AnimalAI experiment described in this paper.
+Appendix G Glossary
 Categorical Random Variable—a discrete random variable that can take on of finite $K$ possible states.
@@
 Factor Graph—bipartite graph representing the factorization of a probability distribution, with one part representing factor nodes and another—random variables.
+Factor size—number of variables connected to the factor. For simplicity, we refer only to the number of previous hidden states, used for prediction.
 Hidden Markov Model (HMM)—statistical model of a stochastic process where state probability depends only on previous state of the process.
 Multi-compartment neuron model—a model of neuron that divides neuron’s connections into groups (segments) of different types (compartments), where each group may be considered as partly independent computational unit and groups of each compartment may affect the neuron’s activity differently.
-Sparse Distributed Representations (SDR)—sparse binary vector in a high-dimensional space, usually formed by k-WTA algorithms.
-Spatial Pooler (SP)—a distributed noise-tolerant online clustering neural network algorithm that converts input binary patterns into SDRs with fixed sparsity while retaining pairwise similarity.
+Oracle—a mythical creature that gives up to our algorithms insightful information about the environment that is usually hidden.
 Successor Representations (SR)—a discounted sum of future [one-hot encoded] observations.

### v4

@@
 Abstract
-This paper presents a novel approach to address the challenge of online temporal memory learning for decision-making under uncertainty in non-stationary, partially observable environments.
-The proposed algorithm, Distributed Hebbian Temporal Memory (DHTM), is based on factor graph formalism and a multicomponent neuron model.
-DHTM aims to capture sequential data relationships and make cumulative predictions about future observations, forming Successor Features (SF).
-Inspired by neurophysiological models of the neocortex, the algorithm utilizes distributed representations, sparse transition matrices, and local Hebbian-like learning rules to overcome the instability and slow learning process of traditional temporal memory algorithms like RNN and HMM.
-Experimental results demonstrate that DHTM outperforms LSTM and a biologically inspired HMM-like algorithm, CSCG, in the case of non-stationary datasets.
-Our findings suggest that DHTM is a promising approach for addressing the challenges of online sequence learning and planning in dynamic environments.
+This paper presents a novel approach to address the challenge of online sequence learning for decision making under uncertainty in non-stationary, partially observable environments.
+The proposed algorithm, Distributed Hebbian Temporal Memory (DHTM), is based on the factor graph formalism and a multi-component neuron model.
+DHTM aims to capture sequential data relationships and make cumulative predictions about future observations, forming Successor Features (SFs).
+Inspired by neurophysiological models of the neocortex, the algorithm uses distributed representations, sparse transition matrices, and local Hebbian-like learning rules to overcome the instability and slow learning of traditional temporal memory algorithms such as RNN and HMM.
+Experimental results show that DHTM outperforms LSTM, RWKV and a biologically inspired HMM-like algorithm, CSCG, on non-stationary data sets.
+Our results suggest that DHTM is a promising approach to address the challenges of online sequence learning and planning in dynamic environments.
 1 Introduction
-Modelling sequential data is one of the essential tasks in artificial intelligence as it has many applications, including decision-making and learning world models [1], natural language processing [2], conversational AI [3], time series analysis [4], and video and music generation [5].
-One of the classical approaches to modelling sequential data is to form a representation that stores and condenses the most relevant information about a sequence, and to find a general transformation rule of this information through the dimension of time [6, 7, 8].
-We refer to the class of algorithms that use this approach as Temporal Memory (TM) algorithms because they essentially model the cognitive ability of complex living organisms to remember the past experience and make future predictions based on that memory [9, 10, 11, 12].
-This paper addresses the problem of online sequence learning for planning and decision-making under uncertainty, which can be formalized as reinforcement learning (RL) for a partially observable Markov decision process (POMDP) [13, 14].
+Modelling sequential data is one of the essential tasks in artificial intelligence as it has many applications, including decision-making and learning world models (Moerland et al., 2023), natural language processing (Min et al., 2021), conversational AI (Dwivedi et al., 2023), time series analysis (Eraslan et al., 2019), and video and music generation (Ji et al., 2020).
+One of the classical approaches to modelling sequential data is to form a representation that stores and condenses the most relevant information about a sequence, and to find a general transformation rule of this information through the dimension of time (Lipton et al., 2015; Harshvardhan et al., 2020; Mathys et al., 2011).
+We refer to the class of algorithms that use this approach as temporal memory (TM) algorithms because they essentially model the cognitive ability of complex living organisms to remember the past experience and make future predictions based on that memory (Hochreiter & Schmidhuber, 1997; Friston et al., 2016; 2018; Parr & Friston, 2017).
+Temporal memories differ in their ability to capture data dependencies and the extent to which they can generalise their predictions to unseen sequences.
+There are two extremes known in psychology as cognitive maps Tolman (1948), which extract commonalities or schemas generalising experience, and episodic memory Tulving et al.
+(1972), which stores fragments of experience separately.
+They both can be viewed as temporal memories, however, the former is slower and require a lot of data to built up a useful schema, according to contemporary computational models based on HMM (George et al., 2021) and Transformers (Dedieu et al., 2024), while the latter is instant and is usually viewed as a data buffer (Zhang et al., 2021).
+In this paper, we explore advantages of an episodic-like memory in addressing the problem of online sequence learning for planning and decision-making under uncertainty, which can be formalized as reinforcement learning (RL) for a partially observable Markov decision process (POMDP) (Poupart, 2005; Singh et al., 2021).
 Inferring the hidden state in a partially observable environment is a sequence modelling problem, since it requires processing a sequence of observations to obtain enough information about the hidden state.
-One of the most efficient representations of hidden states for discrete POMDPs is the Successor Representation (SR), which disentangles hidden states and goals given by the reward function [15, 16, 17].
-The Successor Features framework is an extension of SR to continuous POMDP, using the same idea of value function decomposition, but instead for features of a hidden state [18, 19].
+One of the most efficient representations of hidden states for discrete POMDPs is the Successor Representation (SR), which disentangles hidden states and goals given by the reward function (Dayan, 1993; Samuel J.
+Gershman, 2018; Millidge & Buckley, 2022).
+The Successor Features framework is an extension of SR to continuous POMDP, using the same idea of value function decomposition, but instead for features of a hidden state (Barreto et al., 2017; Touati et al., 2022).
 Temporal memory (TM) algorithms can be used to make cumulative predictions about future states and their features, forming SR or SF.
-This work shows that the proposed algorithm, Distributed Hebbian Temporal Memory (DHTM), can dynamically form SFs for navigation tasks in Gridworld and AnimalAI [20] environments.
-The most prominent TM algorithms, such as a Recurrent Neural Network (RNN) [21] or LSTM [9], use backpropagation to capture data relationships, which is known to be unstable due to recurrent nonlinear derivatives.
+This work shows that the proposed algorithm, Distributed Hebbian Temporal Memory (DHTM), can dynamically form SFs for navigation tasks in Gridworld and AnimalAI (Fountas et al., 2020) environments.
+The most prominent TM algorithms, such as a Recurrent Neural Network (RNN) (Boden, 2002) or LSTM (Hochreiter & Schmidhuber, 1997), use backpropagation to capture data relationships, which is known to be unstable due to recurrent nonlinear derivatives.
 They also require complete data sequences to be available during training.
-Although the gradient vanishing problem can be partially circumvented in a way that Receptance Weighted Key Value (RWKV) [22] or Linear Recurrent Unit (LRU) [23] models do, the problem of online learning, i.e.
+Although the gradient vanishing problem can be partially circumvented in a way that Receptance Weighted Key Value (RWKV) (Peng et al., 2023) or Linear Recurrent Unit (LRU) (Orvieto et al., 2023) models do, the problem of online learning, i.e.
 learning on non-stationary data sets, is still a viable issue.
-Unlike graphical models such as HMM [24], RNN models and their descendants also lack a probabilistic theoretical foundation, which is advantageous for modelling sequences captured from stochastic environments [25, 26].
-There is little research on TM models that can be used in fully online adaptive systems interacting with partially observable stochastic environments with access to only one sequence data point at a time, a common case in reinforcement learning [27].
-We propose a Distributed Hebbian Temporal Memory (DHTM) algorithm based on factor graph formalism and a multi-compartment neuron model inspired by the HTM neuron [28].
-The resulting graphical structure of our model is similar to that of the factorial HMM [29].
+Unlike graphical models such as HMM (Eddy, 2004), RNN models and their descendants also lack a probabilistic theoretical foundation, which is advantageous for modelling sequences captured from stochastic environments (Salaün et al., 2019; Zhao et al., 2020).
+There is little research on TM models that can be used in fully online adaptive systems interacting with partially observable stochastic environments with access to only one sequence data point at a time, a common case in reinforcement learning (Jahromi et al., 2022).
+We propose a Distributed Hebbian Temporal Memory (DHTM) algorithm based on factor graph formalism and a multi-compartment neuron model inspired by the HTM neuron (Hawkins & Ahmad, 2016).
+Factor graph is a flexible graphical representation of probabilistic distributions, which is commonly used in combination with sum-product algorithm (Loeliger, 2004).
+We found this representation particularly convenient in our attempt of probabilistic interpretation of the HTM framework.
+The resulting graphical structure of our model is similar to that of the factorial HMM (Ghahramani & Jordan, 1995).
 An important feature of our model is that the transition matrices for each factor are stored as different components (segments) of artificial neurons, which makes computation very efficient in the case of sparse transition matrices.
-Our TM forms sequence representations entirely online and uses only local Hebbian-like learning rules [30, 31, 32], which avoid the drawbacks of gradient methods and make the learning process much faster than gradient methods.
+Our TM forms sequence representations entirely online and uses only local Hebbian-like learning rules (Hebb, 2005; Churchland & Sejnowski, 1992; Lillicrap et al., 2020), which avoid the drawbacks of gradient methods and make the learning process much more sample efficient than gradient methods.
 This comes at the expense of hidden state generalization, since DHTM’s hidden state uniquely encodes each observation sequence.
 Different trajectories leading to the same position in the partially observable Gridworld can result in different hidden states inferred by DHTM.
-In essence, this makes DHTM a model of episodic memory [33], or an adaptive trajectory buffer, with built-in machinery for making multistep future predictions.
-In this light, we also show that an agent forming SF with DHTM can be interpreted as a kind of episodic control [34, 35, 36, 37].
-The proposed TM is tested as a world model for an RL agent architecture navigating in a Gridworld environment and a more challenging AnimalAI testbed [38].
-The results demonstrate that our algorithm outperforms classical LSTM and a biologically inspired HMM-like world model CSCG [39] in changing environments and is in order of magnitude more sample efficient than Dreamer V3 [40].
+In essence, this makes DHTM a model of episodic memory, or an adaptive trajectory buffer, with built-in machinery for making multistep future predictions.
+In this light, we also show that an agent forming SF with DHTM can be interpreted as a kind of episodic control (EC) (Lengyel & Dayan, 2007; Pritzel et al., 2017; Ritter et al., 2018; Emukpere et al., 2023).
+Though our approach is similar to EC, we are not aware of other works, demonstrating its application to partially observable environments.
+There are other temporal memories that, similarly to DHTM, use Hebbian rules and learn online, like Asymmetric Hopfield Network (AHN) (Chaudhry et al., 2023), Temporal Predictive Coding (tPC) (Tang et al., 2023) and Hierarchical Temporal Memory (Hawkins & Ahmad, 2016).
+Compared to DHTM, AHN is a first-order memory.
+That is, it is not able to disentangle identical observations that correspond to different underlying environmental states.
+Therefore, it is not suitable for partially observable environments.
+In contrast, tPC has a hidden layer that, potentially, makes it a high-order memory.
+Despite that it has been tested only in MDP environments (Tang et al., 2024) so far, this method paves a promising direction, which we aim to compare with in future work.
+Finally, the algorithm that inspired our work, HTM, is also a Hebbian high-order temporal memory, which, opposed to DHTM, doesn’t allow for probabilistic predictions due to its all-or-nothing computations.
+The proposed TM is tested as an episodic memory for an RL agent architecture navigating in a Gridworld environment and a more challenging AnimalAI testbed (Crosby et al., 2020).
+The results demonstrate that our algorithm outperforms classical LSTM, RWKV and a biologically inspired HMM-like world model CSCG (George et al., 2021) in changing environments and is in order of magnitude more sample efficient than Dreamer V3 (Hafner et al., 2023).
 Another advantage of our algorithm is that it can be implemented on neuromorphic processors using only local learning rules.
@@
 2 Background
-This section provides basic information about some of the concepts necessary to follow the paper.
 2.1 Reinforcement Learning
-This paper considers decision-making in a partially observable environment, usually formalized as a partially observable decision process [13].
-A POMDP is defined as a tuple $\mathcal{M}=(S,A,P,R,O,D,\gamma)$, where $S$—state space, $A$—action space, $P(s,a,s`)=Pr(s`\ |\ s,a)$—transition function, $R(s)$–reward function, O—observation space, $D(a,s`,o)=Pr(o\ |\ a,s`)$—sensor model and $\gamma\in[1,0)$—discount factor, given a transition $s,a\to s`$, where $s\in S$, $a\in A$, $o\in O$.
-If $S,A,O$ are finite, $P,D$ can be viewed as real-valued matrices, otherwise, they are conditional density functions.
-Here we consider deterministic rewards that depend only on the current state, i.e.
-$R(s):S\to\mathbb{R}$.
+This paper considers decision-making in a partially observable environment (Poupart, 2005).
+The environment is defined by its state transition function $P(s,a,s^{\prime})=Pr(s^{\prime}\ |\ s,a)$, determining its dynamics, where $s^{\prime},s\in S$ is called state space, $a\in A$—action space.
+Partial observability means that an agent has access to observations $o\in O$, which only partially inform about the environment’s state $s$ at each moment.
 […]
@@
 2.2 Successor Representation
-Successor representations are such representations of hidden states from which we can linearly infer the state value given the reward function [15].
-We assume that the observation and state spaces are discrete.
+Successor representations are such representations of hidden states from which we can linearly infer the state value given the reward function (Dayan, 1993).
+For a discrete state space, such representation can be derived just by rearranging sums in the equation for value function:
 $V(h_{t}=i)=\mathrm{E}[\sum^{\infty}_{l=0}\gamma^{l}R_{t+l+1}\ |\ h_{t}=i]=\sum_{j}\mathrm{SR}_{ij}R_{j},$ (1)
@@
 $R_{j}$ is a reward for observing the state $j$.
-An extension of SR for large or continuous state spaces that we use is Successor Features [18], which is based on a similar idea of value function decomposition, but in a feature space instead (see details in Appendix A).
+An extension of SR for large or continuous state spaces that we use is Successor Features (Barreto et al., 2017), which is based on a similar idea of value function decomposition, but in a feature space instead (see details in Appendix A).
 It is easy to show that, similarly to the Q-function, both SR and SF can be obtained using dynamic programming.
@@
 2.3 Hidden Markov Model
-A partially observable Markov process can be approximated by a Hidden Markov Model (HMM) with hidden state space $H$ and observation space $O$.
-$O$ is the same as in $\mathcal{M}$, but $H$ is generally not equal to $S$.
-Variables $H_{t}$ represent an unobservable (hidden) approximate state of the environment that evolves over time, and observable variables $O_{t}$ represent observations that depend on the same time-step state $H_{t}$, and $h_{t},o_{t}$ are corresponding values of these random variables.
-For simplicity, we assume that actions are fully observable and that information about them is contained in $H_{t}$ variables.
-For the process of length $T$ with state values $h_{1:T}=(h_{1},\ldots,h_{T})$ and $o_{1:T}=(o_{1},\ldots,o_{T})$, the Markov property yields the following factorization of the generative model:
+A partially observable environment can be approximated by a Hidden Markov Model (HMM) with hidden state space $H$ and observation space $O$.
+Generally, $H$ is not equal to environemnts state space $S$ since the environemnt’s true structure is usually unknown.
+For simplicity, we assume that actions are fully observable and that information about them is contained in hidden state.
+For the process of length $T$ with state values $h_{1:T}=(h_{1},\ldots,h_{T})$ and observations $o_{1:T}=(o_{1},\ldots,o_{T})$, the Markov property yields the following factorization of the generative model:
 $p(o_{1:T},h_{1:T})=p(h_{1})\prod_{t=2}^{T}p(h_{t}|h_{t-1})\prod_{t=1}^{T}p(o_{t}|h_{t}).$ (2)
-In the case of a discrete hidden state variable, a time-independent stochastic transition matrix can be learned with the Baum-Welch algorithm [41], a variant of the expectation maximization (EM) algorithm.
-To compute the statistics for the expectation step, it uses the forward-backward algorithm, which is a special case of the sum-product algorithm [42].
-Since the transition matrix grows quadratically with the size of the hidden state space, it may become infeasible to train the HMM in large state spaces.
-Since it is possible that not all hidden states are dependent on each other, the Factorial HMM can mitigate this problem by partitioning the hidden state spaces into smaller independent subspaces, providing an efficient way to handle distributed representations.
-We base DHTM on ideas from the Baum-Welch algorithm and the Factorial HMM, while adapting them for online learning.
-2.4 Episodic Control
-Episodic Control (EC) is a general computational approach to storing and reusing an agent’s experience, inspired by the notion of episodic memory [33] from psychology.
-[…]
-In this paper, we refer to EC to better explain the principle of the proposed algorithm for solving RL tasks.
+In the case of a discrete hidden state variable, a time-independent stochastic transition matrix can be learned with the Baum-Welch algorithm (Baum et al., 1970), a variant of the expectation maximization (EM) algorithm.
+To compute the statistics for the expectation step, it uses the forward-backward algorithm, which is a special case of the sum-product algorithm (Kschischang et al., 2001).
+Factor graph is a convenient graphical representation (Loeliger, 2004) of probabilistic models like HMM (see Fig.1A for an example).
+Formally, a factor graph is a bipartite graph, where one part consists of factors (squares), which, replacing conditional probabilities, define dependencies between random variables (circles) constituting the other part.
+Such representation also allows extension of forward-backward algorithm to sum-product algorithm.
+As shown in section 3.1, we base DHTM on ideas from the sum-product algorithm and the Factorial HMM, while adapting them for online learning.
 3 Distributed Hebbian Temporal Memory
@@
 Distributed Hebbian Temporal Memory is based on the sum-product belief propagation algorithm in a factor graph (see Figure 1A).
-Analogous to Factorial HMM [44], we divide the hidden space $H$ into subspaces $H^{k}$.
+Analogous to Factorial HMM (Ghahramani & Jordan, 1997), we divide the hidden space $H$ into subspaces $H^{k}$.
 There are four sets of random variables (RV) in the model: $H^{i}_{t-1}$—latent variables representing hidden states from the previous time step (context), $H^{k}_{t}$—latent variables for the current time step, and $\Phi^{k}_{t}$–feature variables.
@@
 RV state values are denoted by corresponding lowercase letters: $h^{i}_{t-1}$, $h^{k}_{t}$, $\varphi^{k}_{t}$, $o^{lm}_{t}$.
-Each variable $\Phi^{k}_{t}$ is considered independent and has a separate graphical model for increased computational efficiency.
+In this work, each variable $\Phi^{k}_{t}$ is considered independent and has a separate graphical model for increased computational efficiency.
 However, in practice, hidden variables of the same time step are statistically interdependent.
-We introduce their interdependence through a segment computation trick that goes beyond the standard sum-product algorithm (see equation 6).
+We introduce their interdependence through a segment computation trick that goes beyond the standard sum-product algorithm (see Appendix B for details).
 The model also has three types of factors: $M^{i}_{t-1}$—messages from previous time steps, $F^{k}_{c}$—context factor (generalized transition matrix), $F^{k}_{e}$–emission factor.
@@
 We will discuss only the upper block of the graph in Figure 1A, which is DHTM itself.
-The lower block—an encoder—is described in Appendix C.
+The lower block—an encoder—is described in Appendix D.
 The only requirement for the encoder is that its output should be represented as states of categorical variables (features) for the current observation.
@@
 $p(h_{t}^{k})\propto\sum_{\Omega_{k}}\prod_{i\in\omega_{k}}M_{t-1}^{i}(h_{t-1}^{i})F_{c}^{k}(h_{t}^{k},\Omega_{k}),$ (3)
-where $\Omega_{k}=\{h^{i}_{t-1}:i\in\omega_{k}\}$, $\omega_{k}={i_{1},\dots,i_{n}}$—set of previous time step RV indexes included in $F_{c}^{k}$ factor, $(n+1)$—factor size.
+where summation is over all possible hidden RV states in $\Omega_{k}=\{h^{i}_{t-1}:i\in\omega_{k}\}$, $\omega_{k}={i_{1},\dots,i_{n}}$—set of previous time step RV indexes included in $F_{c}^{k}$ factor, $(n+1)$—factor size.
 3.2 Neural Implementation
-For computational purposes, we translate the factor graph to the Hebbian learning neural network architecture (for a biological interpretation of the model, see Appendix B).
-As shown in Figure 1B, each RV can be viewed as a set of spiking neurons representing the states of the RV, i.e.
-$p(h^{k}_{t})=p(c^{j}_{t}=1)$, where $j$–is the index of a neuron corresponding to the state $h^{k}_{t}$.
-Cell activity is binary $c^{j}_{t}\in\{0,1\}$ (spike/no spike), and the probability can be interpreted as a spike rate.
-Factors $F^{k}_{c}$ and $M^{i}_{t-1}$ can be represented as vectors, where elements are factor values for all possible combinations of RV states included in the factor.
-Let’s denote the elements of the vectors as $f_{l}$ and $m_{u}$ respectively, where $l$ corresponds to a particular combination of $k$ state values $h^{k}_{t},h^{i_{1}}_{t-1},\dots,h^{i_{n_{l}}}_{t-1}$, and $u$ indexes all neurons representing states of RVs from previous time steps.
-Inspired by biological neural networks, we group the connections of a neuron into dendritic segments.
-A segment acts as an independent computational unit that detects a particular input pattern (a contextual state) defined by its receptive field.
-In our model, a segment connects the factor value $f_{l}$ and the excitation $E_{l}$ induced by the segment $l$ to the cell to which it is connected.
-The segment is active, i.e.
-$s_{l}=1$ if all its presynaptic cells are active, otherwise $s_{l}=0$.
-Computationally, a segment transmits its factor value $f_{l}$ to the cell to which it is attached if the context matches the corresponding state combination.
-We can now rewrite equation 3 as the following:
-$p(h_{t}^{k})\propto\sum_{l\in\mathrm{seg}(j)}L_{l}f^{k}_{l},$ (4)
-where $L_{l}=\prod_{u\in\mathrm{rec}(l)}m_{u}$ is segment’s likelihood as long as messages are normalized to probability distributions, $\mathrm{seg}(j)$—indexes of segments that are attached to cell $j$, $\mathrm{rec}(l)$—indexes of presynaptic cells that constitute receptive field of a segment with index $l$.
+For computational purposes, we translate the factor graph shown in Figure 1A to the Hebbian learning neural network architecture shown in Figure 1B (for a biological interpretation of the model, see Appendix C).
+Each RV can be viewed as a set of neurons representing the states of the RV, i.e.
+$p(h^{k}_{t})=p(c^{j}_{t}=1)$, where $j$–is the index of a neuron corresponding to the state $h^{k}_{t}$ and $c^{j}_{t}\in\{0,1\}$—cell’s activity.
+Since RVs are categorical, only one cell per variable can be active at a time and their activity is defined through sampling of RV’s distribution.
+Since factors $F^{k}_{c}$ represent dependencies between variables, in neural implementation we view them as collections of connections between cell groups corresponding to RVs that $F^{k}_{c}$ takes as arguments.
+Each factor’s entry maps a particular combination of states $h^{k}_{t},h^{i_{1}}_{t-1},\dots,h^{i_{n_{l}}}_{t-1}$ to a factor value $f_{l}$.
+We found it useful to represent each such entry as a separate computational unit, which we call segment (alluding to dendritic segments of a biological neuron).
+As shown in Figure 1B, segment $l$ connects its presynaptic cells $h^{i_{1}}_{t-1},\dots,h^{i_{n_{l}}}_{t-1}$ (one per each RV), constituting its receptive field, to the cell it excites $h^{k}_{t}$.
+The factor $f_{l}$ value can be thought of as a segment’s excitation efficiency, i.e.
+how strong its presynaptic pattern influences the cell’s activation.
+Reformulation of belief propagation step in equation 3 in terms of segmental computations results in the following equations:
+$p(h^{k}_{t})=\underset{j\in\mathrm{cells}[H^{k}_{t}]}{\mathrm{softmax}}(\underset{l\in seg(j)}{\mathrm{max}}(E_{l})),$ (4)
+$E_{l}=\log f_{l}+w_{l}\log\frac{1}{n_{l}}\sum_{u\in\mathrm{rec}(l)}w_{ul}m_{u}+\sum_{u\in\mathrm{rec}(l)}(1-w_{ul})\log m_{u},$ (5)
+where $\mathrm{cells}[H^{k}_{t}]$—indexes of cells corresponding to states of the variable $H^{k}_{t}$, $seg(j)$—all segments for cell $j$, $w_{ul}$—synapse efficacy or neuron’s specificity for segment, such that ${w_{ul}=p(s_{l}=1|c^{u}_{t-1}=1)}$, $w_{l}=\frac{1}{n_{l}}\sum_{u}w_{ul}$—average synapse efficacy for segment $l$, and $n_{l}$-number of cells in segment’s receptive field $rec(l)$.
+$m_{u}$ is presynaptic cell’s activation probability $p(c^{u}_{t-1}=1)$, which carries messages from the previous time-step and corresponds to entry in $M^{i}_{t-1}$.
+For detailed derivation and rationale behind these equations, see Appendix B.
 Initially, cells have no segments.
 As learning progresses, new non-zero connections, grouped into segments, are created.
-In equation 4, we benefit from having sparse factor value vectors because their complexity depends linearly on the number of non-zero components.
+In equation 5, we benefit from having sparse factors because their complexity depends linearly on the number of non-zero entries or segments.
 And that’s usually the case in our model due to the one-step Monte Carlo learning and the special form of the emission factors $F^{k}_{e}$:
@@
 where $\mathbb{I}$—indicator function, $\mathrm{col}(\varphi^{k}_{t})$ is a set of hidden states connected to the feature state $\varphi^{k}_{t}$ forming a column.
-The shape of the emission factor is inspired by the presumed columnar structure of the neocortex and has been shown to induce a sparse transition matrix in HMM [39].
-The segment likelihood $L_{l}$ resulting from the sum-product algorithm is computed as if the presynaptic cells were independent.
-However, this isn’t usually the case for sparse factors.
-To account for their interdependence, we substitute the following equation for segment log-likelihood:
-$\log L_{l}$ $=w_{l}\log\frac{1}{n_{l}}\sum_{u\in\mathrm{rec}(l)}w_{ul}m_{u}+\sum_{u\in\mathrm{rec}(l)}(1-w_{ul})\log m_{u},$ (6)
-where $w_{ul}$—synapse efficacy or neuron specificity for segment, such that ${w_{ul}=p(s_{l}=1|c^{u}_{t-1}=1)}$, $w_{l}=\frac{1}{n_{l}}\sum_{u}w_{ul}$—average synapse efficacy for segment $l$, and $n_{l}$-number of cells in segment’s receptive field.
-The idea behind the formula is to approximate between two extreme cases.
-The first is ${p(s_{l}=1|c^{u}_{t-1}=1)}\to 1$ for all $u$, which means that all cells in the receptive field are dependent and part of a cluster, i.e.
-they fire together.
-In this case, ${p(s_{l})=m_{u}}$ for all $u$, but we also reduce the prediction variance by averaging across different $u$.
-In the opposite case, ${p(s_{l}=1|c^{u}_{t-1}=1)\to 0}$ for all $u$, presynaptic cells don’t cluster and the segment activation probability is just a product of the activation probability of each cell.
-The resulting equation for belief propagation in DHTM is the following:
-$p(h^{k}_{t})=p(c^{j}_{t}=1)=\underset{j\in\mathrm{cells}[H^{k}_{t}]}{\mathrm{softmax}}(\underset{l\in seg(j)}{\mathrm{max}}(E_{l})),$ (7)
-where ${E_{l}=\log f_{l}+\log L_{l}}$, $\mathrm{cells}[H^{k}_{t}]$—indexes of cells representing states for the variable $H^{k}_{t}$.
-Here we also approximate the logarithmic sum with the $\mathrm{max}$ operation, inspired by the neurophysiological model of segment aggregation by cells [45].
-The next step after computing the $p(h^{k}_{t})$ distribution parameters is to incorporate information about the current feature states ${p(h^{k}_{t}\ |\ \varphi^{k}_{t})\propto p(h^{k}_{t})\mathbb{I}[h^{k}_{t}\in\mathrm{col}(\varphi^{k}_{t})]}$.
+The shape of the emission factor is inspired by the presumed columnar structure of the neocortex and has been shown to induce a sparse transition matrix in HMM (George et al., 2021).
+The next step after computing the $p(h^{k}_{t})$ distribution is to incorporate information about the current feature states ${p(h^{k}_{t}\ |\ \varphi^{k}_{t})\propto p(h^{k}_{t})\mathbb{I}[h^{k}_{t}\in\mathrm{col}(\varphi^{k}_{t})]}$.
 After that, the learning step is executed.
@@
 First, $h^{i}_{t-1}$ and $h^{k}_{t}$ are sampled from their posterior distributions: ${p(h^{i}_{t-1}\ |\ \varphi^{i}_{t-1})\propto M^{i}_{t-1}}$ and ${p(h^{k}_{t}\ |\ \varphi^{k}_{t})}$ respectively.
-Then $f_{l}$ is updated according to the activity $s_{l}$ of the segment and the activity $c^{j}_{t}$ of its cell, so that $f_{l}$ is proportional to several coincidences ${s_{l}=c^{j}_{t}=1}$ in the recent past:
+Cells that correspond to sampled states become active.
+For cell $j$ corresponding to $h^{k}_{t}$ we check if there are segments matching the pattern of sampled states (active cells) from the previous time-step $\{h^{i}_{t-1}\}_{i}$.
+If there are no such segments, the new segment connecting sampled cells is created.
+For all newly created segments and segments with receptive fields matching active cells from the previous time-step, its activity $s_{l}$ is set to $1$.
+Then $f_{l}$ is updated according to the activity of the segment $s_{l}$ and the activity of its cell $c^{j}_{t}$, so that $f_{l}$ is proportional to several coincidences ${s_{l}=c^{j}_{t}=1}$ in the recent past (exponential moving average of coincidence):
 $\Delta f_{l}=\alpha(c^{j}_{t}-f_{l})s_{l},$ (7)
 where $\alpha\in[0,1)$ is the segment’s learning rate.
-That is, $f_{l}$ is trained to be equal to the average cell activation frequency when the segment is active.
-It’s similar to Baum-Welch’s update rule [41] for the transition matrix in HMM, which in effect counts transitions from one state to another, but in our case, the previous state (context) is represented by a group of RVs.
+That is, $f_{l}$ is trained to be equal to the cell’s average activation frequency when the segment is active.
+It’s similar to Baum-Welch’s update rule (Baum et al., 1970) for the transition matrix in HMM, which, in effect, counts transitions from one state to another, but in our case, the previous state (context) is represented by a group of RVs.
 Weights $w_{ul}$ are also updated by the Hebbian rule to reflect the specificity of a presynaptic $u$ for activating a segment $l$.
 That is, they are targeted to represent the probability ${p(s_{l}=1\ |\ c^{u}_{t-1}=1)}$ that segment $s_{l}$ is active, given that cell $u$ was active at the previous time step.
-We could learn this by counting activation matches and mismatches.
-But in our algorithm, it is approximated as an exponential moving average of the frequency activation of segment $s_{l}$, given $c^{u}_{t-1}=1$:
+In our algorithm, it is approximated as an exponential moving average of the frequency activation of segment $s_{l}$, given $c^{u}_{t-1}=1$:
 $\Delta w_{ul}=\beta\cdot\mathbb{I}[c^{u}_{t-1}=1]\cdot(\mathbb{I}[s_{l}=1]-w_{ul}),$ (8)
@@
 3:     action $\leftarrow$ initial_action
-4:     while (not terminal) (steps $<$ max_steps) do
+4:     while (not terminal) and (steps $<$ max_steps) do
 5:        obs, reward $\leftarrow$ STEP()
@@
 In the OBSERVE() routine, memory learns to predict next feature states as described in section 3.
-An agent learns associations to feature states and rewards in the REINFORCE() function:
+An agent learns associations to map feature states and rewards in the REINFORCE() function:
 $r^{k}_{i}\leftarrow r^{k}_{i}+\alpha\mathbb{I}[\varphi^{k}_{t}=i](R_{t}-r^{k}_{i})$ (9)
@@
 The planning horizon $T$ is determined dynamically from the distribution of $\Phi^{k}_{t+l}$.
-If the distribution of $\Phi^{k}_{t+l}$ is close to uniform or states with positive reward have high probability, the prediction cycle is stopped.
+For each prediction step $l$, we check whether the distribution of $\Phi^{k}_{t+l}$ is close to uniform (according to Kullback-Leibler distance) or whether any of states with positive reward has probability above a threshold; in that case, the prediction cycle is stopped.
 The action value is determined by combining the corresponding SF and feature state rewards as shown in equation 11.
@@
 DTHM performs best when its hidden state space is large.
-Since the hidden state for unpredictable observations is formed randomly, it’s important that there are no accidental state collisions when remembering transitions.
+Since the hidden state for unpredicted observations is formed randomly, it’s important that there are no accidental state collisions when remembering transitions.
 That is, DHTM tries to remember new sequences without interfering with old memories.
@@
 Episodic Control is based on a similar idea, and to investigate its difference from DHTM, we implemented an Episodic Control agent, a simplified counterpart of DHTM, reflecting its core property that presumably helps it solve RL tasks.
+Episodic Control (EC) is a general computational approach to storing and reusing an agent’s experience, inspired by the notion of episodic memory from psychology (Tulving et al., 1972).
+The core of EC is a dictionary $\mathcal{D}_{a}=(K_{a},V_{a})$ that stores key-value pairs for each discrete action $a\in A$.
 In our EC version, $K_{a}$ is an array of hidden states $h_{t}$ and $V_{a}$ is an array of corresponding subsequent hidden states $h_{t+1}$.
@@
 That is, $\mathcal{D}_{a}$ represents the transition matrix of a deterministic HMM for a feature variable $\Phi_{t}$.
-The hidden state $h_{t}$ and the feature state $\varphi_{t}$ have the same dependence as in DHTM (equation 5), but here we assume infinite cells per column.
+The hidden state $h_{t}$ and the feature state $\varphi_{t}$ have the same dependence as in DHTM (equation 6), but in EC we assume that there are infinite hidden states per feature state.
 To predict the next state, we just have to look up $\mathcal{D}_{a}$ for the current state $h_{t}$, but if there is no entry for $h_{t}$ or the prediction does not match the observed state $\varphi_{t+1}$, then the new state is formed.
@@
 3.4.
-However, since the resulting transition matrix is sparse, it’s more efficient to consider $\mathcal{D}_{a}$ as a graph representation and to perform SF formation as a breadth-first search (see Appendix D.3 for details).
+However, since the resulting transition matrix is sparse, it’s more efficient to consider $\mathcal{D}_{a}$ as a graph representation and to perform SF formation as a breadth-first search (see Appendix E.3 for details).
 4 Experiments
@@
 This section shows how different memory models affect the adaptability of an RL agent.
-For each setup, the results are averaged over five experiments with different seed values, and the standard deviation is shown.
-See Appendix D for details on the baselines, their model parameters, and training regimes.
-See Appendix E for a more detailed description of the environments and setups.
-The source code of the experiments and algorithms is at https://anonymous.4open.science/r/dhtm-FA0E.
+For each setup, the results are averaged over five experiments with different seed values, and one standard deviation is shown everywhere, except for boxplots.
+See Appendix E for details on the baselines, their model parameters, and training regimes.
+See Appendix F for a more detailed description of the environments and setups [^1] The source code of the experiments and algorithms is at https://github.com/Cognitive-AI-Systems/him-agent/tree/iclr25..
 4.1 Gridworld
-Figure 2: POMDP 5x5 Gridworld foraging task using different types of memory: DHTM, CSCG, LSTM.
-CSCG and LSTM are trained in two modes.
+Figure 2: POMDP 5x5 Gridworld foraging task using different types of memory: DHTM, CSCG, LSTM, RWKV.
+All agents, except for DHTM, are trained in two modes.
 The first mode is with oracle, which tells the agent to reinitialize memory and clear the trajectory buffer during environmental change.
@@
 The first test was conducted in partially observable Gridworld in a 5x5 alternating maze with a stationary end state and the agent’s starting position.
-The agent can observe the colour of the cell it is in and the colour of an obstacle it encounters, while being punished in addition to the constant step penalty.
-Since the observation can be represented directly as a categorical variable with the number of states equal to the number of possible colours, no encoder is needed.
-Each episode starts with the agent in the same position and ends when the agent enters the terminal state (i.e., when the reward is collected) or when the maximum number of action steps is reached.
+The agent can observe only the colour of the cell it is in and the colour of an obstacle it encounters, while being punished in addition to the constant step penalty.
+Each episode ends when the reward is collected or when the maximum number of action steps is reached.
 In each episode, we measure the number of action steps required for the agent to reach the goal.
 Each trial lasts $1000$ episodes and consists of two phases.
-In the first phase, which lasts $300$ episodes, there are no obstacles on the shortest path to the terminal state.
-In the second phase, the previously optimal path is blocked by a wall, so the agent has to take a detour to get the reward.
+In the first phase, there are no obstacles on the shortest path to the terminal state.
+In the second phase, after $300$ episodes, the previously optimal path is blocked by a wall, so the agent has to take a detour to get the reward.
 The experiment aims to show the ability of a memory to relearn the structure of the maze in the test time.
@@
 Otherwise, the agent will continue to hit the wall and fail to reach the terminal state.
-In this experiment, we compare our memory model DHTM with CSCG [39] and LSTM [9] within the same agent architecture that uses memory to form successor features and evaluate actions (see Sec.
+In this experiment, we compare our memory model DHTM with CSCG (George et al., 2021), LSTM (Hochreiter & Schmidhuber, 1997) and RWKV (Peng et al., 2023) within the same agent architecture that uses memory to form successor features and evaluate actions (see Sec.
 3.4).
-CSCG and LSTM are trained in two modes.
+All memories, except DHTM, are trained in two modes.
 The first mode is with an oracle that tells the agent to reinitialize memory and clear the trajectory buffer just in time for the environmental change.
@@
 The results in Figure 2 show that CSCG effectively can’t adapt to the environmental change in the online regime, but only if it is retrained from scratch on new trajectories.
-And, apparently, LSTM is only able to learn a predictive model for a small horizon, since it performs significantly worse than other baselines in the second phase, even with oracle enabled.
-In contrast, DHTM and EC are able to adapt online without explicit memory reset.
+And, apparently, LSTM and RWKV are only able to learn a predictive model for a small horizon, since they perform significantly worse than other baselines in the second phase, even with oracle enabled.
+In contrast, DHTM and EC are able to adapt online even without explicit memory’s trainable parameters reinitialisation.
 Another important observation is that the EC agent performs on par with DHTM, suggesting that its mechanism can explain most of DHTM’s performance in this task.
+However, see Appendix H.2 for experiments that show their difference in scalability.
 4.2 AnimalAI
-Figure 3: Amount of action steps the agent performs to reach the goal in a 10x10-meter room in AnimalAI environment using DHTM, CSCG, LSTM and EC dictionary.
+Figure 3: Amount of action steps the agent performs to reach the goal in a 10x10-meter room in AnimalAI environment using DHTM, CSCG, LSTM, RWKV and EC dictionary.
 The agent and the goal have the same initial positions throughout the episodes.
@@
 […]
-We test the agent using two different encoders: K-Means and Categorical VAE [46].
+We test DHTM and EC agents using two different encoders: K-Means and Categorical VAE (Jang et al., 2017).
 Both encoders are pre-trained on RGB images collected with a uniform strategy during $2000$ episodes in the same room used for agent testing.
-K-Means represents an image as a single categorical variable with $50$ states, and Categorical VAE is configured for distributed representation with five variables with $50$ states each (see Appendix C for details).
+K-Means represents an image as a single categorical variable with $50$ states, and Categorical VAE is configured for distributed representation with five variables with $50$ states each (see Appendix D for details).
+In case of EC agent, since it doesn’t support distributed representations, we map each unique VAE encoding to one feature state using a simple dictionary.
+Since the resulting number of states for such feature variable is large (typically more than $1000$), it’s infeasible to use the same mapping for CSCG.
+We also didn’t find much difference for LSTM and RWKV using K-Means or VAE, so we only keep the results for VAE encoder.
 […]
@@
 The factor size of 1 corresponds to the graph of the Factorial HMM, with parallel HMMs.
-It can be seen from Figure 3B that in our case of categorical VAE, apparently knowing only the previous state of the variable is not enough to reliably predict rewarding features, since the agent doesn’t converge to the shortest path.
+It can be seen from Figure 3B that in our case of categorical VAE, knowing only the previous state of the variable is not enough to reliably predict rewarding features, since the agent doesn’t converge to the shortest path.
 We also tested the variant (factor 1*) where each feature variable is represented by three hidden states, to ensure that the result for factor size $1$ is not due to DHTM’s hidden state collisions.
 DHTM with a factor that takes into account more than one previous time-step hidden variable (excluding the action variable) performs significantly better.
-For diagrams depicting factor graphs used in these experiments, see Appendix D.4.
+For diagrams depicting factor graphs used in these experiments, see Appendix E.4.
+For further experiments showing the importance of distributed encoding in DHTM, see Appendix H.3.
 5 Conclusion
@@
 Potentially, it extends the application of our algorithm to higher state spaces, exploiting their factorizations for more efficient computations and to reduce the number of stored parameters.
-Our results suggest that DHTM can be viewed as an adaptive trajectory buffer for a slower generalizing algorithm, such as a Neural State Space Model-based [47].
+Our results suggest that DHTM can be viewed as an adaptive trajectory buffer for a slower generalizing algorithm, such as a Neural State Space Model (Jiang et al., 2023).
 Through prediction error learning, DHTM can selectively discard and replace obsolete transitions in observed trajectories without disrupting relevant experience.
@@
 where $\mathrm{SF}^{k}_{j}(h_{t})=\sum^{\infty}_{l=0}\gamma^{l}p(\varphi^{k}_{t+l+1}=j\ |\ h_{t})$, $h_{t}=(h^{1}_{t},...,h^{n}_{t})$—hidden state vector of variables $\{H^{k}_{t}\}_{k}$.
-Appendix B Biological Interpretation
+Appendix B Details of Neuronal Implementation
+As shown in Figure 1B, each RV can be viewed as a set of neurons representing the states of the RV, i.e.
+$p(h^{k}_{t})=p(c^{j}_{t}=1)$, where $j$–is the index of a neuron corresponding to the state $h^{k}_{t}$.
+Cell activity is binary $c^{j}_{t}\in\{0,1\}$, and the probability can be interpreted as a spike rate.
+Factors $F^{k}_{c}$ and $M^{i}_{t-1}$ can be represented as vectors, where elements are factor values for all possible combinations of RV states included in the factor.
+Let’s denote the elements of the vectors as $f_{l}$ and $m_{u}$ respectively, where $l$ corresponds to a particular combination of state values $h^{k}_{t},h^{i_{1}}_{t-1},\dots,h^{i_{n_{l}}}_{t-1}$, and $u$ indexes all neurons representing states of RVs from previous time steps.
+Inspired by biological neural networks, we group the connections of a neuron into dendritic segments.
+A segment acts as an independent computational unit that detects a particular input pattern (a contextual state) defined by its receptive field.
+In our model, a segment connects the factor value $f_{l}$ and the excitation $E_{l}$ induced by the segment $l$ to the cell to which it is connected.
+The segment is active, i.e.
+$s_{l}=1$ if all its presynaptic cells are active, otherwise $s_{l}=0$.
+Computationally, a segment transmits its factor value $f_{l}$ to the cell to which it is attached if the context matches the corresponding state combination.
+We can now rewrite equation 3 as the following:
+$p(h_{t}^{k})\propto\sum_{l\in\mathrm{seg}(j)}L_{l}f^{k}_{l},$ (14)
+where $L_{l}=\prod_{u\in\mathrm{rec}(l)}m_{u}$ is segment’s likelihood as long as messages are normalized to probability distributions, $\mathrm{seg}(j)$—indexes of segments that are attached to cell $j$, $\mathrm{rec}(l)$—indexes of presynaptic cells that constitute receptive field of a segment with index $l$.
+Initially, cells have no segments.
+As learning progresses, new non-zero connections, grouped into segments, are created.
+In equation 14, we benefit from having sparse factor value vectors because their complexity depends linearly on the number of non-zero components.
+And that’s usually the case in our model due to the one-step Monte Carlo learning and the special form of the emission factors $F^{k}_{e}$:
+$F^{k}_{e}(h^{k}_{t},\varphi^{k}_{t})=\mathbb{I}[h^{k}_{t}\in\mathrm{col}(\varphi^{k}_{t})],$ (15)
+where $\mathbb{I}$—indicator function, $\mathrm{col}(\varphi^{k}_{t})$ is a set of hidden states connected to the feature state $\varphi^{k}_{t}$ forming a column.
+The shape of the emission factor is inspired by the presumed columnar structure of the neocortex and has been shown to induce a sparse transition matrix in HMM (George et al., 2021).
+The segment likelihood $L_{l}$ resulting from the sum-product algorithm is computed as if the presynaptic cells were independent.
+However, this isn’t usually the case for sparse factors.
+To account for their interdependence, we substitute the following equation for segment log-likelihood:
+$\log L_{l}$ $=w_{l}\log\frac{1}{n_{l}}\sum_{u\in\mathrm{rec}(l)}w_{ul}m_{u}+\sum_{u\in\mathrm{rec}(l)}(1-w_{ul})\log m_{u},$ (16)
+where $w_{ul}$—synapse efficacy or neuron specificity for segment, such that ${w_{ul}=p(s_{l}=1|c^{u}_{t-1}=1)}$, $w_{l}=\frac{1}{n_{l}}\sum_{u}w_{ul}$—average synapse efficacy for segment $l$, and $n_{l}$-number of cells in segment’s receptive field $rec(l)$.
+The idea behind the formula is to approximate between two extreme cases.
+The first is ${p(s_{l}=1|c^{u}_{t-1}=1)}\to 1$ for all $u$, which means that all cells in the receptive field are dependent and part of a cluster, i.e.
+they fire together.
+In this case, ${p(s_{l})=m_{u}}$ for all $u$, but we also reduce the prediction variance by averaging across different $u$.
+In the opposite case, ${p(s_{l}=1|c^{u}_{t-1}=1)\to 0}$ for all $u$, presynaptic cells don’t cluster and the segment activation probability is just a product of the activation probability of each cell.
+The resulting equation for belief propagation in DHTM is the following:
+$p(h^{k}_{t})=p(c^{j}_{t}=1)=\underset{j\in\mathrm{cells}[H^{k}_{t}]}{\mathrm{softmax}}(\underset{l\in seg(j)}{\mathrm{max}}(E_{l})),$ (17)
+where ${E_{l}=\log f_{l}+\log L_{l}}$, $\mathrm{cells}[H^{k}_{t}]$—indexes of cells representing states for the variable $H^{k}_{t}$.
+Here we also approximate the logarithmic sum with the $\mathrm{max}$ operation, inspired by the neurophysiological model of segment aggregation by cells (Stuart & Spruston, 2015).
+The next step after computing the $p(h^{k}_{t})$ distribution parameters is to incorporate information about the current feature states ${p(h^{k}_{t}\ |\ \varphi^{k}_{t})\propto p(h^{k}_{t})\mathbb{I}[h^{k}_{t}\in\mathrm{col}(\varphi^{k}_{t})]}$.
+After that, the learning step is executed.
+The loop closing step of our TM algorithm is to assign the posterior of the current step ${p(h^{k}_{t}\ |\ \varphi^{k}_{t})}$ to $M^{i}_{t-1}$.
+Appendix C Biological Interpretation
 Neural implementation of the DHTM is inspired by neocortical neural networks (see Fig.
@@
 Hidden variables $H^{k}$ may be considered as populations of excitatory pyramidal neurons in cortical layer L2/3 of somatosensory areas, with lateral inhibition modelled as $\mathrm{softmax}$ function.
-Staiger and Petersen [48] showed that neurons in this layer are responsible for temporal context formation.
+Staiger & Petersen (2021) showed that neurons in this layer are responsible for temporal context formation.
 The neuronal activity at timestep $t$ can be thought to carry messages $M^{k}_{t-1}$.
 Messages are propagated through synapses of dendritic segments, which correspond to factors $F^{k}_{c}$.
-Dendritic segments of biological neurons are known to be coincidence detectors of their synaptic input [45].
-We use the notion of dendritic segment to sparsely represent context factors $F^{k}_{c}$, as each factor value corresponds to a particular combination of states (or active cells).
+Dendritic segments of biological neurons are known to be coincidence detectors of their synaptic input (Stuart & Spruston, 2015).
+That is, each segment can be thought of as a detector of a particular combination of active cells (states in our interpretation).
+If we associate a segment’s parameter with factor value, a collection of dendritic segments represent a mapping from state space to factor values, representing context factor $F_{c}$.
+Our computational model of neuron and its dendrites is based on HTM neuron model (Hawkins & Ahmad, 2016).
+In this model, each segment has a receptive field and activates when a number of active cells in this receptive field reaches a threshold.
+The neuron is active if any of its segments are active.
+We use the same activation rule, but the threshold is fixed to the receptive field size.
+Our algorithm of segment’s growth on demand is also similar to HTM’s.
+In principle, DHTM can be considired as a probabilistic interpretation of HTM, allowing belief propagation.
+Similarly to HTM framework, our temporal memory can be interpreted as a model of cortical layers.
 Feature variables $\Phi^{k}_{t}$ may be considered to represent cells of a granular layer (L4), as they are known to be the main hub for sensory excitation for L2/3.
-L2/3 cells that have common sensory input from the layer L4 are modelled as columns for particular feature states $\mathrm{col(\varphi^{k}_{t})}$ [49].
+L2/3 cells that have common sensory input from the layer L4 are modelled as columns for particular feature states $\mathrm{col(\varphi^{k}_{t})}$ (Mountcastle, 1997).
 Figure 4: Biological view of the neural implementation of the DHTM.
@@
 Spike frequency of a neuron reflects state probability $p(h^{k}_{t})$ of a variable.
-Appendix C Encoding Observations
+Appendix D Encoding Observations
 Since DHTM is designed to process categorical random variables, we need a categorical encoder for RGB images to solve AnimalAI tasks.
@@
 […]
-Appendix D Implementation details
+Appendix E Implementation details
 As mentioned in Section 3.4, we incorporate DHTM as a part of an RL agent, which has a memory model and a feature reward function.
 All tested memory models share the same pipeline—conditioned on actions, they learn sequences of categorically encoded observations and are used to form SFs.
-In all experiments, the agent uses softmax exploration with temperature equal to $0.04$, except for EC agent we set softmax temperature to $0.01$, since for higher temperatures its policy is unstable.
+In all experiments in Section 4.1, the agent uses softmax exploration with temperature equal to $0.04$, except for EC agent we set softmax temperature to $0.01$, since for higher temperatures its policy is unstable.
 For SF formation the agent uses $\gamma=0.8$ and $\gamma=0.9$ for Gridworld and AnimalAI experiments respectively.
@@
 We use probability thresholds equal to $0.05$ and $0.01$ for Gridworld and AnimalAI respectively.
-D.1 LSTM
-LSTM baseline was implemented with a single LSTMCell from PyTorch library [52].
+In scaling Gridworld experiments (Appendix H.2), we increase maximum planning horizon up to $100$ steps, and we also had to decrease agent’s temperature to $0.02$ for DHTM and to $0.002$ for EC agent.
+Discount factor is $0.7$ for DHTM and $0.9$ for EC.
+We found that, in 10x10 maze, DHTM’s strategy becomes more stable if we increase the positive reward threshold up to $0.4$.
+EC agent converges faster if we increase its softmax temperature, however, it converges to less optimal trajectory on average.
+In the noise tolerance AnimalAI experiment (Appendix H.3), for DHTM we use the same set of agent parameters, except we set the goal probability threshold to $0.5$.
+We also set the goal detection threshold equal to the mean positive reward since we found that noise variable states mostly get reward assignment slightly above zero, in contrast to signal variables, in which only several states get positive reward assignment.
+E.1 LSTM and RWKV
+LSTM baseline was implemented with a single LSTMCell from PyTorch library (Paszke et al., 2019).
 It is supported by an additional symexp-layer to encode input before passing it to the LSTM cell and a symexp-layer to decode the LSTM cell’s output from the LSTM’s hidden state back to the input representation, where symexp activation function, $\mathrm{symexp}(x)=\mathrm{sign}(x)e^{\lvert x\rvert-1}$, is a reverse of symlog function: $\mathrm{symlog}=\mathrm{sign}(x)\log{(\lvert x\rvert+1})$.
-For both Gridworld and AnimalAI tasks, LSTM with a hidden state size of 100 is trained on a buffer of $1000$ trajectories with learning rate $0.002$ on $50$ randomly sampled batches of size $50$ every $10$ episodes.
-We also incorporated some notion of random variables in hidden states by splitting the hidden state of the tested LSTM into groups.
+The similar way we implemented RWKV baseline: a single RWKV layer supported by single-layer encoder and decoder.
+Current public RWKV implementation is a fast evolving framework (Bo, 2021), and for the increased performance it is tightly bound to the offline batch training common for the transformer architectures.
+In our case we needed a so-called sequential mode for online learning similar to LSTM.
+Thus, we adapted another public implementation mentioned in the official documentation (RWKV in 150 lines of code).
+For both Gridworld and AnimalAI tasks, LSTM and RWKV with a hidden state size of 100 are trained on a buffer of $1000$ trajectories with learning rate $0.002$ on $50$ randomly sampled batches of size $50$ every $10$ episodes.
+We also incorporated some notion of random variables in hidden states by splitting the hidden state of the tested RNNs into groups.
 In all experiments the hidden state represents $10$ categorical variables with $10$ states.
-That is, LSTM is forced to learn $10$ categorical distributions with multi-cross-entropy loss to explain the observed sequences, which is a somewhat close to the multi-categorical hidden state representation used in DreamerV2/V3 [40].
+That is, RNN is forced to learn $10$ categorical distributions with multi-cross-entropy loss to explain the observed sequences, which is a somewhat close to the multi-categorical hidden state representation used in Dreamer V2/V3 (Hafner et al., 2023).
 The idea of using symexp activation function, mentioned above, is inspired by Dreamer too, and is used to remedy the problem of learning extreme probability values.
@@
 Thus, symexp function makes it faster to reach target values in log space.
-D.2 CSCG
+E.2 CSCG
 CSCG baseline was implemented using code from the repository accompanying the paper (https://github.com/vicariousinc/naturecomm_cscg).
 CSCG was trained on buffers of $500$ action steps with $1000$ EM steps until convergence.
-We iteratively calculated the exponential moving average of transition matrices obtained for different batches with a smoothing coefficient $\alpha=0.05$ as described in the paper [53].
+We iteratively calculated the exponential moving average of transition matrices obtained for different batches with a smoothing coefficient $\alpha=0.05$ as described in the paper (Dedieu et al., 2019).
 This smoothed transition matrix was used as an initialization for the next training batch and for inference.
 The first batch of observations is gathered using the uniform policy, and later batches are gathered using the current best policy.
-In so-called oracle mode, the trajectory buffer is continuously growing, and the transition matrix is trained from scratch every $500$ steps.
-The buffer is discarded at the environmental change and the transition matrix is reinitialised.
-D.3 Episodic Control
+In so-called oracle mode, the transition matrix is trained from scratch every $500$ steps. The buffer is discarded at the environmental change and the transition matrix is reinitialised.
+E.3 Episodic Control
 Here, we provide an algorithm for SF formation using EC dictionary $\mathcal{D}_{a}$ (see Algorithm 3).
-In contrast to other memory algorithms, EC dictionary doesn’t estimate state transition probabilities, but it stores connections between them in all-or-nothing faction.
+In contrast to other memory algorithms, EC dictionary doesn’t estimate state transition probabilities, but it stores connections between them in all-or-nothing fation.
 Therefore, on each time step, all predicted observations are assumed equally probable.
@@
 16:  end for
-D.4 DHTM
+E.4 DHTM
 In the Gridworld task and AnimalAI with K-Means encoder, the factor-graph for DHTM consists of three hidden variables $H$ with $40$ states per one feature state.
@@
 takes into account tree hidden variables from the previous time-step).
-Appendix E Experimental Setups
-E.1 Gridworld
+Appendix F Experimental Setups
+F.1 Gridworld
 We use a simple homebrew Gridworld implementation.
@@
 Samples of setups used in our experiments are depicted in Figure 6.
-Here, the red colour corresponds to obstacles and the light gray colour to the reward.
+Additional setups for scaling experiments are shown in Figure 7.
+Here, the red colour corresponds to obstacles (negative numbers) and the light gray colour to the reward (number 4).
+Letters A and G correspond to the agent’s initial position and goal, correspondingly.
 Floor colours are resampled for each trial, but we ensure that there are no more than five floor colours, including the terminal state.
@@
 Figure 6: Schematic view of POMDP Gridworld setups.
-The left picture shows the initial setup of the environment, with the reward delivered by the light gray square.
-The right picture represents the environment after the reward is blocked by obstacles (red squares).
-Agent’s initial position in the lower left corner.
-E.2 AnimalAI
-AnimalAI is a testbed inspired by experiments with animals [38].
+The left picture shows the initial setup of the environment, with the reward delivered by the light gray square (number 4).
+The right picture represents the environment after the reward is blocked by obstacles (red squares or negative numbers).
+The letter A depicts the agent’s initial position and the letter G—the goal.
+Figure 7: Schematic view of POMDP Gridworld setups for scale experiments. From left to right: 5x5, 7x7, 10x10. The letter A depicts the agent’s initial position and the letter G—the goal.
+F.2 AnimalAI
+AnimalAI is a testbed inspired by experiments with animals (Crosby et al., 2020).
 The environment consists of 3D area surrounded by a wall and many different objects that can be placed using a configuration file, including walls, food, ramps, trees, movable obstacles, and more.
 We tested our agent in a 10x10-meter room surrounded by walls (see Fig.
-7).
+8).
 Each timestep, the agent gets an RGB image of a first-person view of the environment and is able to influence its behaviour by choosing one of three actions: turn left, turn right, or go forward.
@@
 Figure 8: Sample of the agent’s first-person view of our setup in AnimalAI environment. The green ball represents a goal, a food item to be collected.
-Appendix F Additional Experiments
+Appendix G Capacity
+In this section, we provide basic theoretical analysis of DHTM’s capacity.
+Since the hidden state for unpredicted feature state transition is generated randomly, the main source of errors is accidental collision of hidden states.
+Due to special form of the emission factor (see equation 6), the collision is possible only for the same feature sate.
+Here we analise the capacity of DHTM for factor grpah depicted in Figure 5.
+Let’s assume that the stream of feature states is sampled from uniformly distributed categorical random variable.
+Therefore, the probability to have a collision $p_{c}=\frac{1}{n_{\varphi}n_{v}n_{h}}$, where $n_{v}$—number of hidden variables, $n_{\varphi},n_{h}$ are the number of states for feature and hidden variables correspondingly.
+Therefore, the probability of having no collisions for a $l$ transitions is equal to $(1-p_{c})^{l}$.
+Let’s set a threshold $c\in(0,1)$, which determines our confidence in that the sequence of transitions is collision free.
+Solving the equation $1-c=(1-p_{c})^{l}$ with respect to $l$, we get:
+$l=\frac{\log(1-c)}{\log(1-\frac{1}{n_{\varphi}n_{v}n_{h}})}.$ (18)
+Using the Taylor expansion for $n_{\varphi}n_{v}n_{h}\to+\inf$, we can see that $l$ grows linearly with the total number of states $n_{\varphi}n_{v}n_{h}$.
+Therefore, DHTM’s capacity scales linearly with the number of neurons.
+This, however, can be easily improved if we set additional prior on hidden state activation probability, which would take into account hidden state usage, but we leave it for future research.
+Appendix H Additional Experiments
+H.1 Dreamer
 In order to ground our experiments to state-of-the-art model based algorithms, as an additional baseline, we test the smallest version of Dreamer V3 with 18 million parameters.
 We use the implementation described in Voudouris et al.
-[54], which is based on Hafner et al.
-[40] and adapted for AnimalAI.
+(2023), which is based on Hafner et al.
+(2023) and adapted for AnimalAI.
 […]
 Figure 9: Dreamer V3 (18M) score in AnimalAI experiment described in this paper.
-Appendix G Glossary
+H.2 Scalability
+In this section, we explore DHTM’s scalability properties and compare them to the EC agent.
+In RL experiments, like presented in section 4, the number of segments largely depends on the RL algorithm used, its exploration strategy, and how fast it converges to a stable policy.
+We tested DHTM and EC agents in empty Gridworld of different sizes: 5x5, 7x7 and 10x10, where the agent and the goal are placed in opposite corners (see Fig.
+7).
+As shown in Figure 10A, the DHTM-based agent requires fewer episodes to find the goal; however, the policy becomes less stable as the size of the Gridworld grows.
+As a consequence of faster convergence, DHTM produces fewer segments than the number of EC’s dictionary entries throughout experiments.
+It also can be seen that for both DHTM and EC agents, the number of memory units grows non-linearly with the state space size (see Fig.
+10B).
+Figure 10: Comparison of EC and DHTM-based agents in foraging task in POMDP Gridworld of different sizes: 5x5, 7x7 and 10x10.
+A.
+Dynamic of steps required for the agent to reach the goal.
+B.
+Number of segments grown (or dictionary entries for EC agent) after $300$ episodes with respect to state space size.
+In principle, the number of segments reflects the number of transitions stored, which is proportional to the amount of experience provided to the memory.
+To show that, we tested DHTM and EC agent in 10x10 empty Gridworld using the same memory parameters as for experiments in section 4.1, but under uniform strategy.
+That is, agents freely explored the environment during $100$ episodes and at the end the number of segments for DHTM and dictionary size for EC was recorded for each episode length.
+Figure 11 shows that DHTM’s number of segments and EC’s dictionary size grow linearly with the maximum episode length.
+In other words, since DHTM lacks generalisation abilities it largely serves as a trajectory buffer, in the same sense as a dictionary in EC.
+Figure 11: Comparison of DHTM’s number of segments with EC agent’s dictionary size depending on episode length in Gridworld after uniform exploration during $100$ episodes.
+DHTM’s computational complexity (time and memory) of the prediction step grows linearly with respect to the number of segments.
+Therefore, to use it for huge environments, there is need to prune unnecessary segments or remember only selected episodes.
+This is a common problem among episodic-like memories.
+The current approach to form SF also should be adjusted to be scalable, since it requires maximum planning horizon to be equal to the episode length.
+Among the improved approaches could be to employ n-step TD learning, hierarchical planning or SF caching, to recalculate it less frequently.
+H.3 Noise Tolerance
+In order to show the key difference between DHTM and its simplified EC version, we tested agents in the same AnimalAI setup as described in 4.2, but we add one noise variable to the VAE encoder.
+It’s easy to see that the naive mapping from distributed VAE encoding to one categorical variable used for EC doesn’t work in this case, since even identical observations will likely be encoded differently due to the noise variable.
+In contrast, since DHTM is able to handle distributed representations directly (as the name implies), it still can use non-noise variables independently to learn transitions.
+In this experiment, we vary DHTM’s factor size (see Fig.
+12).
+Since factors formed completely random, probability to include the noisy variable increases with factor size and this consequently worsen the performance of the agent.
+If the factor size is set to the total number of variables, DHTM becomes effectively unfactorised, similarly to its EC counterpart.
+Figure 12: Number of action steps the agent takes to reach the goal in a 10x10-meter room in the AnimalAI environment, using DHTM and EC models with a distributed Categorical VAE encoder that includes one additional noisy variable (six variables in total).
+DHTM’s factor size is varied, determining the number of variables sampled for each factor.
+Appendix I Glossary
 Categorical Random Variable—a discrete random variable that can take on of finite $K$ possible states.